### v3

# Abdoul Kader Kaboré

Abstract
Unit testing is essential for software reliability, yet manual test creation is time-consuming and often neglected. Although search-based software testing improves efficiency, it produces tests with poor readability and maintainability. Although LLMs show promise for test generation, existing research lacks comprehensive evaluation across execution-driven assessment, reasoning-based prompting, and real-world testing scenarios.
This study presents the first large-scale empirical evaluation of LLM-generated unit tests at the class level, systematically analyzing four state-of-the-art models—GPT-3.5, GPT-4, Mistral 7B, and Mixtral 8x7B—against EvoSuite across 216,300 test cases from Defects4J, SF110, and CMD (a dataset mitigating LLM training data leakage). We evaluate five prompting techniques—Zero-Shot Learning (ZSL), Few-Shot Learning (FSL), Chain-of-Thought (CoT), Tree-of-Thought (ToT), and Guided Tree-of-Thought (GToT)—assessing syntactic correctness, compilability, hallucination-driven failures, readability, code coverage metrics, and fault detection capabilities.
Our findings challenge prior claims that in-context learning is ineffective for test generation in code-specialized LLMs. Reasoning-based prompting—particularly GToT—significantly enhances test reliability, compilability, and structural adherence in general-purpose LLMs. However, hallucination-driven failures remain a persistent challenge, manifesting as non-existent symbol references, incorrect API calls, and fabricated dependencies, resulting in high compilation failure rates (up to 86%). Execution-based classification and mutation testing reveal that many failing tests stem from hallucinated dependencies, limiting effective fault detection.
Keywords:
Automatic Test Generation · Unit Tests · Large Language Models · Prompt Engineering · Empirical Evaluation
1 Introduction
Unit testing is fundamental to software quality, yet inadequate testing remains pervasive—with only 48% of developers regularly writing tests (Daka and Fraser, 2014). While testing ensures correctness by validating components before integration (Beck, 2000), the labor-intensive nature of test creation leads to significant coverage gaps in practice. Furthermore, when tests do exist, their poor structure often creates brittle, incomprehensible test suites that hamper maintenance (Panichella et al., 2020; Almasi et al., 2017; Grano et al., 2018).
Traditional automated test generation approaches have attempted to address these challenges, but each faces significant limitations. Search-Based Software Testing (SBST), particularly tools like EvoSuite (Fraser and Arcuri, 2011), optimizes for code coverage but produces tests that are difficult to read (Almasi et al., 2017), contain unintuitive naming (Daka et al., 2015), and rely on generic assertions (Panichella et al., 2020). Most critically, these high-coverage tests often fail to effectively detect faults (Pinto and Vergilio, 2010)—the fundamental purpose of testing. Alternative approaches like fuzzing (Zeller et al., 2019), random testing (Pacheco and Ernst, 2007), and symbolic execution (Sen et al., 2005) each introduce their own trade-offs between coverage, precision, and scalability.
Large Language Models (LLMs) (Brown et al., 2020; Fan et al., 2023) have emerged as a promising alternative for test generation, potentially addressing the fundamental limitations of traditional approaches. Initial studies suggest LLM-generated tests feature more natural naming conventions, human-like assertions, and better readability than SBST alternatives (Siddiq et al., 2024a; Yuan et al., 2023; Tang et al., 2024). Moreover, LLMs can generate both functional and exploratory tests that go beyond mere structural coverage (Wang et al., 2024)—a flexibility that aligns with comprehensive testing practices.
Despite their promise, LLM-driven test generation remains largely unproven for production use. Current evaluations focus primarily on syntactic correctness rather than practical integration and execution outcomes. Our research identifies and addresses five critical gaps that must be resolved before LLM-based testing can be reliably deployed:
  • Impact of Prompt Engineering on Test Reliability and Usability: Prior work (Tang et al., 2024; Siddiq et al., 2024a) has focused on minimal prompting strategies, leaving structured reasoning approaches (CoT, ToT, GToT) underexplored. We evaluate their effectiveness in improving test structure, compilability, and consistency.
  • Structural Correctness & Extractability of LLM-Generated Tests: Prior work examined compilability (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024) but not structural correctness and extractability. We apply MSR and CSR metrics (Macedo et al., 2024) to assess whether tests are valid JUnit cases and can be seamlessly integrated into testing pipelines.
  • Execution Reliability and Fault Detection: SBST tests maximize coverage but often lack valid assertions (Fraser and Arcuri, 2011; Shamshiri et al., 2015). LLM tests show high failure rates (30.4% for ChatGPT) (Yang et al., 2024), reducing executable tests for bug detection. ChatGPT detects bugs in only 21% of cases versus 26% for EvoSuite (Yang et al., 2024). We address gaps in prior work by categorizing test errors and applying mutation testing with Pitest to assess assertion validity and defect detection.
  • Readability, Maintainability & Test Smells: LLM tests are considered more readable than SBST outputs (Yuan et al., 2023; Tang et al., 2024), but readability doesn’t ensure maintainability. Test smells reduce long-term usability (Panichella et al., 2020). We assess maintainability using readability models (Scalabrino et al., 2018; Sergeyuk et al., 2024), complexity metrics, and automated test smell detection, comparing LLMs with EvoSuite.
  • Generalizability Across Diverse Codebases: Prior studies on LLM-generated tests rely on limited benchmarks (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024), raising concerns about generalizability. We address this by evaluating three datasets: Defects4J, featuring real-world buggy software; SF110, containing Java programs with existing test cases; and CMD, a curated dataset excluding projects released before May 2023 to mitigate LLM training data leakage.
This paper: We systematically evaluate LLM-generated test suites and prompt engineering impact. Building on prior comparisons (Tang et al., 2024; Siddiq et al., 2024a; Yang et al., 2024), surveys (Wang et al., 2024), and AI risk studies (Sallou et al., 2023), we conduct a large-scale evaluation across four LLMs and five prompting techniques against EvoSuite. Unlike method-level approaches in prior work, our class-level evaluation aligns with real-world testing workflows.
We make the following key contributions:
  1. Systematic Investigation of Prompt Engineering: We expand on previous research (Tang et al., 2024) by testing five structured prompting techniques (ZSL, FSL, CoT, ToT, GToT) with general-purpose LLMs, moving beyond the code-specific models in  Yang et al. (2024). Unlike prior method-level studies (Siddiq et al., 2024a; Yang et al., 2024; Yuan et al., 2024; Chen et al., 2024) our class-level evaluation provides a more realistic assessment by capturing inter-method dependencies. Our results challenge claims (Yang et al., 2024) that reasoning-based prompting has limited impact on test generation for code-specialized LLMs. Our experiments demonstrate GToT consistently enhances test quality—improving structure, reducing hallucinations, and increasing extractability and compilability.
  2. Systematic Structural Correctness & Extractability Analysis: Our analysis reveals that reasoning-based prompting improves test extractability and real-world usability, addressing inconsistencies in LLM test suite generation.
  3. Systematic Execution Outcome & Fault Detection Analysis: We classify test outcomes as Passed, Failed, or Timeout, revealing gaps between syntactic correctness and execution behavior. We apply mutation testing to assess fault detection by introducing synthetic faults and measuring their detection rate. However, most LLM-generated Passed tests proved unsuitable for this evaluation as they primarily targeted interfaces, abstract classes, or trivial methods without engaging actual program logic. This resulted in a 0% mutation score, revealing that while LLMs can generate syntactically correct tests that pass, these tests often lack engagement with core program logic—making them unsuitable for mutation testing and limiting their ability to verify meaningful program behavior.”
  4. Systematic Differentiation Between Readability & Maintainability: We analyze how prompt engineering impacts both readability and long-term maintainability, highlighting areas where LLMs produce cleaner or more problematic test structures.
  5. Systematic Multi-Dataset, Multi-Model Evaluation for Generalizability: We benchmark four LLMs on Defects4J, SF110, and CMD, offering a broader dataset scope than prior studies. CMD mitigates data leakage by including projects released after May 2023.
  6. Publicly Available Benchmark & Replication Package: We provide a benchmark of 216,300 tests across 690 Java classes, with prompt templates, evaluation scripts, and results accessible at: https://anonymous.4open.science/r/LLM4TS-0F76/
The remainder of this paper is structured as follows: Section 2 covers LLM-based test generation, Section 3 outlines our experimental setup, Section 4 presents evaluation results, Section 5 discusses threats to validity, Section 6 reviews related work, Section 7 discusses key findings and implications, and Section 8 concludes the study.
2 Background
In this section, we introduce the fundamental concepts essential for understanding our paper. We cover LLMs for unit test generation along with prompt engineering, which constitutes the central ideas of this study.
2.1 LLM-based Unit Test Generation
Recent research has explored LLMs for automated test generation. Unlike traditional generators, LLMs leverage pre-trained knowledge to generate tests from natural language or code contexts. Schafer et al. (Schäfer et al., 2023) evaluated LLMs’ ability to produce valid test code. Siddiq et al. (Siddiq et al., 2024a) assessed compilability and coverage of ChatGPT-generated tests, revealing execution reliability gaps. Tang et al. (Tang et al., 2024) showed LLMs lag behind SBST tools in fault detection despite producing syntactically valid tests. Studies by Yuan et al. (Yuan et al., 2024) and Chen et al. (Chen et al., 2024) demonstrated prompt structure affects test quality. Yang et al. (Yang et al., 2024) provided cross-LLM evaluation, while Bhatia et al. (Bhatia et al., 2024) found LLM-generated Python tests suffer from incorrect assertions. Enhancement efforts include TestArt (Gu et al., 2024) and CITYWALK (Zhang et al., 2025).
Current works often focus narrowly on static correctness or coverage, lacking assessment of maintainability, runtime behavior, or test smells. Our study addresses these limitations with a holistic framework combining execution-based validation, mutation testing, and maintainability analysis. We compare five prompt engineering techniques, demonstrating how reasoning-driven prompting improves reliability, structure, and usability across three diverse datasets.
Figure 1: A Sample Use of ChatGPT in Unit test Suite generation
2.2 Prompt Engineering
Prompt engineering creates effective prompts that guide LLMs in specific tasks (Amatriain, 2024; Siddiq et al., 2024b; Naveed et al., 2023; Sahoo et al., 2024). This approach is valuable for complex problems where single-line reasoning is insufficient (Amatriain, 2024; Kojima et al., 2022; Wei et al., 2022), allowing LLMs to consider multiple possibilities before reaching conclusions. The significance of prompt engineering is twofold: specific prompts yield precise responses (Si et al., 2022), and well-crafted instructions control output format, reducing post-processing needs (Macedo et al., 2024). Software engineering applications typically use basic approaches like Zero-shot (Kojima et al., 2022) and Few-shot learning (Brown et al., 2020), though recent research has expanded to include Chain-of-Thought (Wei et al., 2022), self-consistency (Wang et al., 2022), and Tree-of-Thoughts (Wang et al., 2022).
Zero-shot learning enables models to generalize without examples but struggles with complexity (Sahoo et al., 2024). Few-shot learning uses examples to improve performance on novel tasks (Brown et al., 2020). Chain-of-Thought (CoT) (Wei et al., 2022) outlines intermediate reasoning steps, while Self-consistency (Wang et al., 2022) uses diverse reasoning paths to select consistent answers. Tree-of-Thoughts (ToT) (Long, 2023) builds on CoT by exploring a reasoning ”tree” with search algorithms. This study applies these techniques to adapt LLMs for unit test generation.
2.3 Match Success Rate (MSR) and Code Extraction Success Rate (CSR)
Match Success Rate (MSR) and Code Extraction Success Rate (CSR), introduced by Macedo et al. (2024), evaluate the reliability of LLM-generated code. MSR measures the proportion of outputs where predefined patterns successfully extract source code, ensuring structural adherence. CSR assesses the proportion of valid test code extracted, filtering out explanatory text or formatting inconsistencies. These metrics are crucial for usability assessment, as LLM outputs often mix code with commentary (Figure 1).
2.4 Readability vs. Understandability
Readability and understandability are distinct code quality aspects. Readability concerns visual clarity (indentation, naming, formatting), while understandability measures cognitive effort needed to comprehend logic (Buse and Weimer, 2009; Oliveira et al., 2022; Grano et al., 2018; Winkler et al., 2024). Readability models evolved from traditional approaches (Buse and Weimer, 2009; Daka et al., 2015; Posnett et al., 2011; Dorn, 2012) to deep learning-based assessments (Mi et al., 2022). Though human evaluations are ideal, their scalability is limited by subjectivity and evaluator fatigue (Kumar et al., 2024; Zhang et al., 2024a; Sun et al., 2024; Sergeyuk et al., 2024).
Scalabrino et al. (2018) implemented a robust readability model integrating syntactic, structural, visual, and textual elements from Java projects. Using 104 features, it improved accuracy by 6.2% over previous methods and better matched human assessments Sergeyuk et al. (2024). Its training on jUnit projects makes it especially suitable for evaluating test code readability. Unit tests present unique challenges—poor readability obscures intent and hinders maintenance, particularly with complex fixtures (Grano et al., 2018; Winkler et al., 2024). Our study evaluates LLM-generated tests for readability, maintainability, and practical usability.
2.5 Test Smells
Unit testing verifies individual components’ functionality while maintaining readability. Test smells indicate poor practices (Palomba et al., 2016a; Peruma et al., 2019) that lead to fragility and tight coupling. Assertion Roulette and Magic Number Tests—documented in TsDetect [^1] https://testsmells.org/pages/testsmellexamples.html—commonly appear in both EvoSuite and LLM-generated tests. In Listing 1, Assertion Roulette shows multiple assertThat() calls without clarifications, obscuring failure sources.
1@MediumTest
2public void testCloneNonBareRepoFromLocalTestServer() throws Exception {
3 Clone cloneOp = new Clone(false, integrationGitServerURIFor(”small-repo.early.git”), helper().newFolder());
4 Repository repo = executeAndWaitFor(cloneOp);
5 assertThat(repo, hasGitObject(”ba1f63e4430bff267d112b1e8afc1d6294db0ccc”));
6 File readmeFile = new File(repo.getWorkTree(), ”README”);
7 assertThat(readmeFile, exists());
8 assertThat(readmeFile, ofLength(12));
9}
Listing 1: Example of Assertion Roulette Test Smell
In Listing 2, Magic Number Test demonstrates unexplained values (15, 30) without context, compromising maintainability.
1@Test
2public void testGetLocalTimeAsCalendar() {
3 Calendar localTime = calc.getLocalTimeAsCalendar(BigDecimal.valueOf(15.5D), Calendar.getInstance());
4 assertEquals(15, localTime.get(Calendar.HOUR_OF_DAY));
5 assertEquals(30, localTime.get(Calendar.MINUTE));
6}
Listing 2: Example of Magic Number Test Test Smell
We use TsDetect (Peruma et al., 2020) to analyze how prompt engineering affects test smell prevalence in LLM-generated tests compared to EvoSuite-generated tests.
3 Experimental Setup
3.1 Overview
Our methodology (Figure 2) evaluates LLM-generated test suites using ZSL, FSL, CoT, ToT, and GToT prompting techniques. To assess prompt engineering and reliability (RQ1, RQ2), we analyze structural correctness and usability using MSR, CSR, syntax correctness, and compilability. For readability and maintainability (RQ3, RQ4), we employ automated metrics including a deep learning model (Scalabrino et al., 2018) that correlates with human assessments (Sergeyuk et al., 2024), alongside complexity metrics and coding standard adherence. For code coverage (RQ5), we compare LLM tests with EvoSuite across line, instruction, and method coverage. We analyze execution outcomes (RQ6) by categorizing tests as Passed, Failed, or Timeout, using mutation testing to assess fault detection. Finally, we evaluate test smell prevalence (RQ7) to determine if reasoning-based prompting improves maintainability.
Figure 2: Overview of the pipeline used to design the experiments.
3.2 Research Questions
This study evaluates LLMs in test generation and the impact of prompt engineering through seven research questions:
RQ1: How do prompt engineering techniques impact the output format and extractability of LLM-generated test suites? We examine how different prompting strategies influence test suite structure and extractability through ❶ Match Success Rate (MSR), measuring detection of test code via predefined patterns, and ❷ Code Extraction Success Rate (CSR), evaluating reliable extraction of valid test code. These metrics determine adherence to JUnit structures and isolation from extraneous content.
RQ2: How does prompt engineering influence the generation of syntactically correct and compilable test suites by LLMs? We assess how prompt engineering helps LLMs generate test suites that: ❶ adhere to syntax rules and ❷ successfully compile, enabling execution without ensuring functionality.
RQ3: Beyond syntactic correctness, how well do LLMs generate human-understandable and maintainable test suites? We evaluate if LLM-generated tests: ❶ convey test logic clearly; ❷ remain adaptable over time; and ❸ integrate seamlessly into workflows, determining if they meet practical development standards.
RQ4: How do LLM-generated test suites compare to EvoSuite and align with developer expectations in terms of readability? We investigate whether LLM tests meet developer expectations by: ❶ enhancing navigation through clear naming and structure; ❷ reducing ambiguity with well-organized presentation; and ❸ supporting collaboration. We analyze how prompting techniques influence readability using an automated model correlated with human evaluation.
RQ5: How Thorough Are LLM-Generated Test Suites in Terms of Code Coverage? We examine code coverage (line, instruction, method) of LLM-generated tests compared to EvoSuite (Fraser and Arcuri, 2011), analyzing how prompting strategies impact exploration of program behavior.
RQ6: How do the execution outcomes of LLM-generated test suites reflect their ability to detect faults and their applicability in real-world scenarios? We analyze test execution outcomes—(1) Passed, (2) Failed, and (3) Timeout—and use mutation testing to assess fault detection capability (Fraser and Arcuri, 2011), uncovering patterns and limitations in LLM-generated tests.
RQ7: How do LLM-generated test suites compare to EvoSuite-generated test suites in terms of test smell prevalence? We compare the absence of test smells in LLM-generated tests against EvoSuite to determine whether prompt engineering produces more maintainable tests with fewer design issues than traditional SBST techniques.
3.3 Large Language Models
Our study focuses on four instruction-tuned decoder-only models known for their enhanced ability to follow instructions compared to base models. These models belong to the OpenAI and MistralAI families and are detailed in Table 1.
Table 1: Summary of LLMs Used in Current Study
Model Name # Parameters (Billions) Context Length (Tokens) Release Date
GPT 3.5-turbo (OpenAI, 2023) 175 4 096 Nov, 2022
GPT 4 (Achiam et al., 2023) >>>175 8 192 Mar, 2023
Mistral 7B (Jiang et al., 2023) 7 4 096 Sep, 2023
Mixtral 8x7B (Jiang et al., 2024) 46.7 8 192 Dec, 2023
GPT-3.5-turbo has a token limit of 4,096, and exceeding this limit can hinder class context understanding, affecting test quality (Tang et al., 2024). To ensure fairness, all models adhere to this limit, excluding inputs that exceed it. Token counts are computed using OpenAI’s tiktoken [^2] https://github.com/openai/tiktoken library.
We select OpenAI models (GPT-3.5, GPT-4) and MistralAI models (Mistral 7B, Mixtral 8x7B) for their relevance and efficiency. OpenAI models ensure comparability with prior LLM-based test generation studies (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024). MistralAI models, underexplored in this domain, offer cost-effective alternatives: Mistral 7B matches Llama 34B’s performance (Jiang et al., 2023), while Mixtral 8x7B, leveraging Sparse Mixture-of-Experts (SMoE), surpasses Llama 2 70B with lower inference costs (Jiang et al., 2024). This selection balances established and emerging models, broadening the scope of LLM-based test generation research.
Furthermore, all models are set to a standard temperature of 0.7. Our selection criteria prioritized models that excel in code generation and possess accessible APIs, making them suitable for both academic research and industrial applications (Wang et al., 2024; Fan et al., 2023).
3.4 Prompt Design
Unit test generation research emphasizes practical and structured approaches (Siddiq et al., 2024a; Yuan et al., 2023; Chen et al., 2023; Tang et al., 2024). To ensure clarity, we design prompts with varying complexity using a standardized template. Each prompt comprises: (1) a Natural Language Description (NLD) specifying the task, and (2) a Source Code placeholder for inserting the Class Under Test (CUT). The following sections detail these components and their alignment with unit testing practices.
  • Natural Language Description: We design Natural Language Descriptions (NLDs) that guide LLMs in test generation, including: (i) a role-playing directive simulating professional testers (Tang et al., 2024; Chen et al., 2023); (ii) a task-specific directive for JUnit 4 test file generation; and (iii) a control statement enforcing structured output with markers. We introduce delimiters (###Test START### and ###Test END###) to separate test code from explanatory text.
  • Source Code: Aligned with prior work on learning-based unit test generation (Tang et al., 2024; Chen et al., 2023), our prompts vary by technique, either including (i) only the class name and source code or (ii) an additional example class with its test suite to provide context and comparative guidance.
This study evaluates five prompt techniques: Zero-Shot Learning (ZSL), Few-Shot Learning (FSL), Chain-of-Thought (CoT), Tree-of-Thought (ToT), and Guided Tree-of-Thought (GToT). ZSL uses minimal context for generalization, while FSL introduces examples to improve structure. For consistency, we use the same medium-sized class across experiments, ensuring token limit compliance for FSL examples. CoT enhances reasoning through step-by-step prompting, while ToT explores multiple reasoning paths before selecting test cases. GToT, our novel approach, simulates collaborative reasoning among three ”testers” who iteratively refine tests, combining CoT’s structured reasoning with ToT’s diversity to improve coverage, correctness, and variability. Figure 3 illustrates CoT and ToT prompts, while Figure 4 shows GToT’s collaborative structure. All prompts are available in our replication package.
[…]
Figure 3: Examples of Chain-of-Thought and Tree-of-Thoughts for test suite generation.
[…]
Figure 4: Example of Guided Tree-of-Thoughts (GToT) for test suite generation.
3.5 The EvoSuite SBST tool
We employ its default settings, including the DynaMOSA algorithm (Panichella et al., 2017), which optimizes branch, method, and line coverage. To mitigate the randomness of genetic algorithms, we run EvoSuite 30 times per class, following (Tang et al., 2024). A three-minute time budget is allocated per test generation, aligning with prior studies (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015), which find this duration sufficient for effective test case generation. We deliberately disable EvoSuite’s post-generation processing, which reduces test smells, to ensure a fair comparison with unoptimized LLM-generated tests. This approach allows for an unbiased evaluation of the baseline capabilities of both methods, highlighting their inherent strengths and limitations.
Hardware
For reference, EvoSuite experiments are conducted on an AMD EPYC 7552 48-Core Processor running at 3.2 GHz with 640 GB of RAM. However, it’s important to note that not all available resources are necessarily utilized during the tests.
3.6 Datasets
To minimize selection bias and ensure generalizability, we employ three diverse datasets: SF110, Defects4J, and a Custom Mini Dataset (CMD). Both SF110 and Defects4J are used to evaluate both EvoSuite and LLM performance. The CMD, containing GitHub projects created after May 2023, is exclusively tested with LLMs due to compatibility issues with EvoSuite’s Java version requirements  [^3] https://github.com/EvoSuite/evosuite/issues/433. This dataset focuses on well-maintained projects from the OceanBase Developer Center (ODC)  [^4] https://github.com/oceanbase/odc and Conductor OSS  [^5] https://github.com/conductor-oss/conductor.
CMD includes modern, actively maintained open-source projects from the OceanBase Developer Center (ODC) and Conductor OSS, offering a complementary perspective to SF110 and Defects4J. While SF110 covers academic and industrial Java programs and Defects4J focuses on real-world bug-fixing scenarios, CMD introduces contemporary, evolving software. Despite its smaller scale, CMD is carefully curated to meet specific constraints for test generation evaluation. Its size is limited by three key factors. First, token constraints are considered, ensuring that each class can accommodate all prompt engineering techniques (ZSL, FSL, CoT, ToT, GToT) without exceeding token limits. Second, the dataset focuses on recent development, including only classes committed after May 2023 to reflect modern Java practices and prevent potential data leakage. Finally, CMD is designed to follow a buggy-fixed pairing approach, similar to Defects4J, where each class is selected with both a buggy and a fixed version, enabling structured and controlled evaluation of test generation performance. ODC (13 releases, 527 stars, 20 contributors) is a database development tool offering SQL development, risk management, and data security. Conductor OSS (11 releases, 15.5k stars, 53 contributors) is a microservices orchestration framework originally developed by Netflix. These modern, actively maintained projects provide a challenging testbed for LLM-generated tests, allowing us to assess generalization beyond the structured environments of SF110 and Defects4J. These projects are selected for their real-world relevance and evolving nature, reflecting modern Java development practices.
The datasets encompass projects with code of varying levels of cyclomatic complexity (detailed in Table 2), ranging from low to high. This diversity helps mitigate potential overfitting on specific complexity profiles and provides valuable insights into the potential of LLMs for real-world automated test case generation.
Table 2: Overview of datasets and their complexity.
Dataset Description Number of Projects Number of Classes Complexity Overview
Low Moderate High
SF110 Derived from DynaMOSA benchmark (Panichella et al., 2017), includes 346 Java classes from 117 projects. 74 182 170 138 120
Defects4J Contains 835 reproducible bugs across 17 open-source projects, focuses on detailed test suites to expose bugs. 16 477 254 177 130
Custom mini Dataset (CMD) Includes ODC, a collaborative project to enhance database functionalities, and Conductor OSS, an open-source project management tool aimed at efficiency and scalability. 2 31 16 9 5
  • 1
    Notice. Due to budgetary reasons, we only run GPT-4 and Mixtral-8x7B on the smallest dataset (CMD). Other models (GPT-3.5-Turbo and Mistral-7B) are run on all datasets.
4 Experimental Results
4.1 [RQ1]: Impact of prompt engineering on the output format and extractability.
[Experiment Goal]: This research examines how prompt engineering affects the format and extractability of LLM-generated test suites, evaluated through Match Success Rate (MSR) for detecting test code and Code Extraction Success Rate (CSR) for extracting structured test code, regardless of its syntactic correctness or compilability.
[Experiment Design]: We evaluate prompt designs using Match Success Rate (MSR) and Code Extraction Success Rate (CSR). MSR measures adherence to predefined structures, while CSR assesses extraction of structured test code excluding commentary. To address LLMs’ tendency to mix code with explanations (Figure 1), we introduced delimiters (###Test START## and ###Test END##) to improve extraction accuracy (Figures 3). Following Macedo et al. (2024), we used regular expressions for extraction. For CSR validity, test suites needed key components like import statements, ensuring completeness rather than isolated snippets. CSR evaluates extractability independent of syntax correctness or compilability. Experiments on SF110, Defects4J, and CMD assessed LLM performance across diverse scenarios.
[Experiment Results]: Our study generated 216,300 files, with 190,619 (88.13%) containing source code. Table 3 summarizes MSR and CSR metrics across datasets and models, with detailed prompt-specific results in Tables 4 and 5.
Dataset-Level Performance. LLMs showed varying MSR and CSR scores across datasets (Table 3). GPT-3.5-turbo achieved the highest MSR (99.64%) and CSR (90.79%) on CMD, indicating strong structured output generation. GPT-4 also performed well, reaching a CSR of 83.59% on CMD. Mistral 7B excelled in SF110 (MSR: 94.00%) and Defects4J (MSR: 91.10%), demonstrating better structural adherence in these datasets. In contrast, Mixtral 8x7B struggled on CMD, with lower MSR and CSR scores.
Table 3: MSR and CSR by LLM and Dataset.
Model Name # of Generated Files # of Files with code MSR CSR
SF110 Defects4J CMD SF110 Defects4J CMD SF110 Defects4J CMD SF110 Defects4J CMD
GPT 3.5-turbo 27 300 71 550 4 650 23 048 60 565 4 633 84.42% 84.65% 99.64% 78.10% 78.34% 90.79%
GPT 41 4 650 4 356 93.68% 83.59%
Mistral 7B 27 300 71 550 4 650 25 664 65 180 4 226 94.00% 91.10% 90.88% 79.97% 80.35% 83.61%
Mixtral 8x7B1 4 650 2 947 63.38% 53.57%
Total/Average (%) 54 600 143 100 18 600 48 712 125 745 16 162 89.21 87.88 86.90 79.04 79.35 77.89
  • 1
    Note that due to budget constraints, we only paid to query GPT-4 and Mixtral 8x7B for CMD.
Finding 1: OpenAI models (GPT-3.5, GPT-4) demonstrate superior MSR and CSR, indicating their ability to generate structured, extractable test suites. Mistral 7B also performs well, whereas Mixtral 8x7B struggles, highlighting model-specific variations.
Impact of Prompting Techniques. GToT consistently achieved the highest MSR across all models and datasets(Table 4), reinforcing its effectiveness in enforcing structural adherence. For GPT-3.5-turbo, it significantly improved MSR, reaching 21.82% on Defects4J, 21.65% on SF110, and 21.94% on CMD. Mistral 7B also benefited from GToT, achieving 20.79% on Defects4J and 21.07% on SF110. Surprisingly, ZSL performed well in certain cases, with GPT-3.5-turbo reaching 21.98% MSR on CMD. In contrast, FSL consistently resulted in lower MSR scores, suggesting that additional examples do not necessarily enhance structural adherence. ToT outperformed CoT, improving MSR by up to 2.85% on SF110, reinforcing the advantage of stepwise reasoning in structuring outputs.
Table 4: MSR by LLM and Prompt Engineering
Model Name Defects4J (%) SF110 (%) CMD (%)
ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT
GPT 3.5-turbo 18.88 16.41 17.99 17.81 21.82 18.03 16.24 17.53 16.93 21.65 21.98 21.98 21.86 21.86 21.94
GPT 4 22.01 21.96 21.68 18.67 18.76
Mistral 7B 19.72 19.31 20.07 20.10 20.79 19.72 19.39 19.63 20.19 21.07 20.04 19.88 20.04 19.83 20.21
Mixtral 8x7B 20.04 19.88 20.04 19.83 20.21
Finding 2: GToT is the most effective prompting strategy for improving adherence to follow prompt instructions (MSR), while ZSL remains a strong baseline, especially for OpenAI models. FSL underperforms, challenging assumptions that more examples lead to better structural adherence.
GToT consistently achieved the highest CSR scores (Table 5), demonstrating its effectiveness in reducing extraneous content. GPT-3.5-turbo reached 21.10% on Defects4J, 19.65% on SF110, and 21.04% on CMD, while Mistral 7B followed a similar trend with 18.95% on CMD. ZSL remained a strong baseline, with GPT-3.5-turbo achieving 21.68% on CMD, indicating that direct instructions can sometimes outperform complex reasoning. FSL performed the worst, suggesting excessive prompt complexity introduces noise. Mixtral 8x7B struggled across all strategies, with its CSR dropping to 13.91% on CMD under GToT.
Table 5: CSR by LLM and Prompt Engineering
Model Name Defects4J (%) SF110 (%) CMD (%)
ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT
GPT 3.5-turbo 18.5 14.13 15.05 17.21 21.10 17.69 14.37 15.0 16.37 19.65 21.68 20.75 20.07 16.37 21.04
GPT 4 21.15 21.06 19.03 16.19 14.55
Mistral 7B 17.87 17.85 18.21 17.76 16.52 18.08 17.99 17.82 17.38 13.81 18.95 18.67 18.74 18.5 17.13
Mixtral 8x7B 10.91 13.77 11.17 9.18 13.91
Finding 3: GToT improves extractability (CSR), reducing unnecessary content. However, ZSL remains competitive for GPT models, while FSL consistently underperforms. Mixtral 8x7B struggles significantly across all prompts.
To address LLMs’ tendency to mix test code with explanatory text, we introduced delimiters (###Test START### and ###Test END###) for structured output. Despite these clear boundaries, no LLM consistently follows the format, with MSR varying between 16% and 22% across models and datasets (Table 5). This inconsistency directly impacts CSR, which ranges from 9% to 22% (Table 5), highlighting persistent challenges in automated test code extraction.
Finding 4: Explicit delimiters improve extraction by structuring LLM outputs, but all LLMs exhibit inconsistencies in following formatting instructions, with some failing more frequently, making automated extraction challenging.
4.2 [RQ2]: Syntactical Correctness and Compilability.
[Experiment Goal]: This research question examines how prompt engineering influences test suite syntactical correctness and compilability. Valid Java grammar and compilation without modifications are essential for automated workflow integration. We aim to identify the most effective strategies for generating valid, executable test suites.
[Experiment Design]: We evaluate syntactical correctness using Javalang  [^6] https://github.com/c2nes/javalang parsing and assess compilability with JVM compilation. Tests are syntactically correct if parsing succeeds and compilable if compilation completes without errors. For non-compilable suites, we categorize compilation errors (e.g., ’Cannot Find Symbol’, ’Package Does Not Exist’) to identify hallucination patterns. We analyze how different prompting techniques affect these error patterns and assess their effectiveness in mitigating specific hallucinations, examining error distribution across strategies. This experiment spans SF110, Defects4J, and CMD datasets to evaluate how approaches influence correctness, compilability, and hallucination mitigation across diverse codebases.
[Experiment Results]:
Dataset-Level Analysis. Table 3 presents a detailed breakdown of the LLM-generated test cases across SF110, Defects4J, and CMD. In total, our study generates 216,300 files, with 190,619 (88.13%) containing source code. Of these, 171,187 are syntactically correct, while 25,681 consist of textual outputs or structurally invalid test cases. These results highlight the challenge LLMs face in consistently generating well-formed unit tests, emphasizing the need for post-generation validation.
Finding 5: Approximately 21% of LLM-generated test suites that contain code fail syntactical correctness checks, requiring manual intervention before integration into real-world workflows.
LLM and Prompting Impact on Syntactical Correctness. Table 6 reveals significant variations in syntactical correctness across models and prompting techniques. GPT-3.5-turbo with GToT achieves the highest rates (23.65% on SF110, 24.54% on Defects4J, 21.06% on CMD), demonstrating structured reasoning’s benefits for code formation. GPT-4 reaches 23.00% on CMD using ZSL, indicating strong encoded syntactical patterns even with minimal prompting, while Mistral 7B performs best with FSL on SF110 (21.15%) and CoT on Defects4J (20.65%), suggesting different models benefit from different prompting approaches. Mixtral 8x7B consistently shows the lowest correctness rates across datasets.
Regarding prompting techniques, GToT consistently outperforms other strategies across models and datasets. Surprisingly, FSL underperforms despite providing examples, while ZSL performs well for GPT models, particularly GPT-3.5-turbo reaching 21.70% on CMD. These results highlight the crucial role of reasoning-based prompting in enhancing syntactical correctness, with different models showing varied responsiveness to specific techniques.
Finding 6: GPT models outperform others in syntactical correctness, with GPT-4 achieving strong results (23.00%) even with ZSL prompting, while Mixtral 8x7B struggles across datasets. GToT consistently improves correctness across all models, while FSL fails to enhance structure learning despite providing examples.
Table 6: Syntactic Correctness Across Datasets and Techniques1
Model Name Syntactically Correct SF110 (%) Syntactically Correct Defects4J (%) Syntactically Correct CMD (%)
ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT
GPT 3.5-turbo 21.30 17.30 18.05 19.70 23.65 21.51 16.43 17.51 20.02 24.54 21.70 20.77 20.09 16.39 21.06
GPT 4 23.00 22.90 20.68 17.60 15.82
Mistral 7B 21.25 21.15 20.95 20.43 16.23 20.26 20.23 20.65 20.13 18.73 20.37 20.29 20.60 20.11 18.62
Mixtral 8x7B 18.85 23.36 18.51 15.58 23.60
Average (%) 21.27 19.22 19.50 20.06 19.94 20.89 18.33 19.08 20.08 21.63 21 21.83 19.97 17.42 19.78
  • 1
    Note that due to budget constraints, we only paid to query GPT-4 and Mixtral 8x7B for CMD.
Compilability Rates Across Datasets and Techniques. Table 7 reveals a significant gap between syntactical correctness and compilability. GPT-3.5-Turbo with GToT achieved the highest rates (9.67% on SF110, 6.76% on Defects4J), yet only 7.2% of generated tests and 12% of syntactically correct ones compiled. On CMD, GPT-3.5-Turbo with CoT surprisingly outperformed GPT-4 (5.92% vs. 2.65%), suggesting execution failures limit GPT-4’s syntax advantage. Mixtral 8x7B showed inconsistent performance (near-zero to 20.16% with GToT), indicating occasional success but poor overall reliability.
Table 7: Compilability Rates of Test Suites Across Datasets and Techniques1
Model Name Compilable Code SF110 (%) Compilable Code Defects4J (%) Compilable Code CMD (%)
ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT ZSL FSL CoT ToT GToT
GPT 3.5-turbo 5.94 4.21 5.46 4.08 9.67 2.95 2.22 3.23 2.26 6.76 2.70 2.44 5.92 2.68 3.46
GPT 4 1.78 2.21 2.65 1.90 2.39
Mistral 7B 1.09 1.04 0.91 1.74 3.37 0.45 0.43 0.27 0.44 1.09 0.00 0.03 0.00 0.03 0.31
Mixtral 8x7B 0.16 0.12 0.08 0.44 20.16
Average (%) 3.52 2.62 3.19 2.91 6.52 1.70 1.32 1.75 1.35 3.93 1.16 1.20 2.16 1.26 1.58
  • 1
    ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-thoughts.
These results contrast with prior work (Siddiq et al., 2024a), which reports higher compilation rates due to method-level test generation. Class-level dependencies and inter-method relationships add complexity, increasing failures as LLMs struggle to resolve imports, hierarchies, and scopes.
Finding 7: The gap between syntactical correctness and compilability remains substantial—only 7.2% of all test suites and 12% of syntactically correct ones successfully compile. Class-level dependencies introduce additional complexity, making compilation significantly harder than the method-level approach used in prior studies.
Compilation Error Analysis and Hallucination Patterns. Table 8 reveals dataset-specific hallucination patterns, with CMD showing highest rates. ’Cannot Find Symbol’ (CFS) errors dominate across models: 86.47% (GPT-3.5-Turbo), 76.99% (GPT-4), 72.41% (Mistral 7B), and 67.91% (Mixtral 8x7B), indicating generation of non-existent symbols. ’Package Does Not Exist’ (PDNE) errors are significant in Mixtral 8x7B (22.35%) and Mistral 7B (23.13%), reflecting incorrect imports.
Table 8: Compilation Errors by Model and Dataset (All Prompting Techniques)
Model Dataset AR ATM CA CFS CR CAM DCD ITB IT MAM MRS MNA PDNE PAI PCR SCI UE UCL USL UET VNI WAP
GPT 3.5-Turbo Defects4J 0.45 0.0 0.83 84.21 0 2.26 0.0 0.04 1.58 0.24 0.01 0 3.53 6.36 0.02 0 0.43 0.01 0 0.04 0 0
SF110 0.45 0.0 9.05 72.40 0 2.53 0.01 0.01 1.83 0.31 0.01 0 1.48 11.67 0.04 0.01 0.19 0 0 0 0 0
CMD 0.21 0 0 86.48 0 0.46 0 0.03 0.51 1.39 0 0.00 7.87 2.96 0.00 0.00 0.08 0.00 0 0.01 0.00 0.00
GPT 4 CMD 0.21 0 0 76.99 0 0.24 0 0.05 1.08 3.03 0 0.00 16.46 1.50 0.11 0.01 0.30 0.01 0 0.00 0.01 0.00
Mistral 7B Defects4J 0.31 0.06 0.34 81.03 0.01 1.27 0.0 0.10 5.11 0.73 0.00 0.01 6.63 3.68 0.04 0.02 0.17 0.43 0.03 0.02 0 0
SF110 0.86 0.06 0.35 82.0 0 1.84 0.0 0.01 5.18 0.97 0.00 0 2.81 5.56 0.02 0.03 0.16 0.15 0.01 0 0 0
CMD 0.29 0 0 72.41 0 0.71 0 0.09 0.86 1.82 0.02 23.13 0.61 0.04 0.00 0.00 0.00 0.00 0 0.00 0.00 0.01
Mixtral 8x7B CMD 0.31 0 0 67.91 0 0.76 0 0.38 1.59 4.77 0.00 22.35 1.81 0.04 0.02 0.05 0.00 0.00 0 0.00 0.00 0.00
  • 1
    AR: Ambiguous Reference, ATM: Array Type Mismatch, CA: Cannot Access, CFS: Cannot Find Symbol, CR: Class Redefinition, CAM: Constructor Argument Mismatch, DCD: Duplicate Class Declaration, ITB: Incompatible Type Bound, IT: Incompatible Types, MAM: Method Argument Mismatch, MRS: Missing Return Statement, MNA: Modifier Not Allowed, PDNE: Package Does Not Exist, PAI: Private Access Issue, PCR: Public Class Redefinition, SCI: Static Context Issue, UE: Unchecked Exception, UCL: Unclosed Character Literal, USL: Unclosed String Literal, UET: Unused Exception in Try, VNI: Variable Not Initialized, WAP: Weaker Access Privileges.
Defects4J reveals encapsulation issues with high CFS (GPT-3.5: 80.57%, Mistral 7B: 82.12%) and PAI errors (GPT-3.5: 7.88%, Mistral 7B: 4.74%), suggesting access permission hallucinations. SF110 shows redundancy problems with prevalent CFS (GPT-3.5: 81.02%, Mistral 7B: 82.42%) and PAI peaking at 11.76% with GPT-3.5 Few-shot. PCR and DCD errors indicate LLMs generate redundant test structures. Our findings confirm  Zhang et al. (2024b)’s conclusions that LLM hallucinations occur due to missing context and poor dependency handling, while extending their work by demonstrating these same patterns specifically in test generation contexts. The prevalence of CFS errors (up to 86.47% in GPT-3.5-Turbo) confirms their observation about fabricated references, while PDNE errors (up to 23.13% in Mistral 7B) extend their findings by revealing the particular challenges of dependency resolution in test suite generation.
Finding 8: LLM hallucinations are a major source of compilation errors. CMD is dominated by hallucinated missing symbols, Defects4J struggles with encapsulation violations, and SF110 suffers from redundant test generation.
Prompt Engineering and Hallucination Mitigation. Figure LABEL:fig:compilation-error-heatmap reveals the distribution of non-zero error types across different machine learning models and prompt engineering techniques, with the y-axis dynamically adapting to display error types specific to each model and prompt configuration. The heatmap (Figure LABEL:fig:compilation-error-heatmap) shows CoT and GToT reduce hallucinations more effectively than ZSL and FSL. CoT lowers ’Private Access Issue’ (PAI) errors to 2.96% (GPT-3.5) and 1.50% (GPT-4) in CMD, improving encapsulation adherence. GToT nearly eliminates errors in Defects4J through structured dependencies. However, hallucinations in symbol resolution and redundant test generation persist despite structured reasoning.
FSL increases PAI errors to 11.31% (GPT-3.5) and 5.64% (Mistral 7B) in SF110, showing persistent access restriction misinterpretations. ’Cannot Find Symbol’ (CFS) errors remain high (86.47% in CMD, 79.16% in SF110), indicating examples don’t resolve symbol hallucinations. ZSL performs worst with CFS errors reaching 86.47% (CMD) and 78.89% (SF110) due to minimal context. Our findings extend  Zhang et al. (2024b, 2025) by demonstrating how reasoning-based prompting like CoT and GToT can reduce certain error types—particularly PAI and CAM—by reinforcing syntactic scaffolding. While reasoning-based prompting reduces errors, challenges in reference resolution, encapsulation, and redundant test generation remain.
Finding 9: Reasoning-based prompting (CoT, GToT) significantly reduces errors, particularly Private Access Issues, whereas Zero-shot Learning exacerbates compilation failures by generating unstructured, non-compilable test cases.
4.3 [RQ3]: Static Analysis for Quality.
[Experiment Goal]: This research question evaluates LLM-generated test suite quality through static analysis of coding standard compliance, complexity, and potential defects. We assess adherence to Java conventions, maintainability based on complexity metrics, and structural flaws to determine practical usability.
[Experiment Design]: We evaluate test quality using three approaches. Checkstyle [^7] https://checkstyle.sourceforge.io assesses adherence to Google Java StyleCheckstyle [^8] https://google.github.io/styleguide/javaguide.html and Sun Code Conventions [^9] https://www.oracle.com/java/technologies/javase/codeconventions-contents.html, measuring formatting consistency and standard violations. PMD [^10] https://pmd.github.io/ analyzes cyclomatic complexity (independent paths) and cognitive complexity (comprehension ease) (Dantas and Maia, 2021). SpotBugs [^11] https://pmd.github.io/ detects potential defects categorized by severity: Scariest (critical issues), Scary (high-risk errors), Troubling (moderate concerns), and Of Concern (minor warnings). We compared prompting techniques (ZSL, FSL, CoT, ToT, GToT) across SF110, Defects4J, and CMD datasets.
[Experiment Results]:
Coding Standard Compliance. LLM-generated test suites show varying coding standard adherence (Figure LABEL:fig:style-heatmap). FSL achieves highest compliance with Google Java Style and Sun Code Conventions, producing fewer violations overall.
For Google Java Style (Figure LABEL:fig:google-heatmap), indentation violations dominate across all prompting techniques, with GToT showing the lowest count (1757.5 vs. 2305.9 for ZSL). However, GToT increases import-related violations (140.6 CustomImportOrder and 93.1 AvoidStarImport violations) compared to FSL (71.3 and 23.4 respectively). For Sun Code Conventions (Figure LABEL:fig:sun-heatmap), MissingJavadocMethod violations are highest with CoT (566.5) and GToT (564.0), while GToT significantly reduces MagicNumber violations (134.8 vs. 323.4 for ZSL). These patterns suggest that reasoning-based prompts (CoT, ToT, GToT) improve code structure but sacrifice documentation and import organization. FSL strikes the best balance, minimizing both structural and documentation violations. These findings partially align with  Tang et al. (2024), who reported high ZSL violations, though our results show FSL and CoT significantly improve compliance while ToT and GToT produce tests with better logical structure but less style adherence.
Finding 10: FSL achieves the highest coding standard compliance, followed by CoT, which also reduces violations and improves adherence to style guidelines. In contrast, GToT enhances logical organization but introduces more violations, reinforcing the trade-off between structured prompting and strict adherence to coding conventions.
Code Complexity Analysis. Table 9 shows most LLM-generated test suites exhibit low complexity across prompting strategies. GToT produces the highest proportion (99.99%) of low-cyclomatic-complexity tests while generating no instances of high cognitive complexity, demonstrating its ability to create structured yet maintainable and understandable test cases.
Table 9: Summary of Complexity Results by Prompt Engineering Techniques1
(a) Cyclomatic Complexity Results
Cyclomatic
Complexity (%)
ZSL FSL CoT ToT GToT
Low (1-4) 99.86 99.50 99.73 99.73 99.99
Moderate (5-7) 0.11 0.39 0.16 0.20 0.01
High (8-10) 0.03 0.11 0.11 0.07 0
Very High (11+) 0 0 0 0 0
(b) Cognitive Complexity Results
Cognitive
Complexity (%)
ZSL FSL CoT ToT GToT
Low (¡5) 99.83 99.76 99.76 99.90 100
Moderate (5-10) 0.11 0.14 0.14 0.03 0
High (11-20) 0.03 0.10 0.11 0.07 0
Very High (21+) 0.03 0 0 0 0
  • 1
    ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
These findings align with Tang et al. (2024), who found ChatGPT tests under ZSL exhibit low complexity. However, our results show reasoning-based prompting (CoT, ToT, GToT) refines logical structure while maintaining this low complexity. While ZSL produces simple tests, it lacks the structured refinement of guided reasoning approaches, highlighting the advantage of stepwise logical improvements in test generation.
Finding 11: GToT produces simpler, more maintainable test suites with the lowest cyclomatic and cognitive complexity. ZSL and CoT also generate structurally sound tests, while FSL primarily improves coding standard compliance without significantly altering complexity.
Static Analysis with SpotBugs. Table 10 reveals 8,461 issues in 20,530 compilable tests (55.58% minor ”Of Concern” warnings). FSL generated the most Scariest issues (61.72%), while GToT produced the most Scary (63.92%) and Troubling (69.10%) defects. ZSL had fewer severe issues but simpler tests. This highlights a key trade-off: structured reasoning improves organization but increases subtle defects. Unlike Tang et al. (2024)’s findings with ZSL, our study shows FSL and GToT’s enhanced structure comes with more severe defects, requiring better balance between structured reasoning and defect mitigation.
Table 10: Bug Priority Levels by Prompt Techniques
Prompt Bug Level GPT-3.5 GPT-4 Mistral7B Mixtral8x7B
SF110 D4J CMD CMD SF110 D4J CMD CMD
ZSL Of Concern 600 212 17 15 129 48 0 2
Scariest 0 0 0 0 129 48 0 0
Scary 0 0 0 0 129 0 0 0
Troubling 0 212 0 15 0 0 0 0
FSL Of Concern 410 177 18 14 135 60 0 0
Scariest 410 177 0 0 135 60 0 0
Scary 0 0 0 0 0 0 0 0
Troubling 0 0 0 14 0 0 0 0
CoT Of Concern 553 219 22 0 125 32 0 0
Scariest 0 0 0 0 0 32 0 0
Scary 0 0 0 0 125 0 0 0
Troubling 0 0 0 0 125 0 0 0
ToT Of Concern 20 21 3 7 212 63 1 0
Scariest 0 1 0 0 212 63 0 0
Scary 1 1 0 0 212 0 0 0
Troubling 0 2 0 1 0 0 0 0
GToT Of Concern 825 242 20 19 347 131 0 4
Scariest 0 0 0 0 0 0 0 0
Scary 825 0 0 0 0 0 0 0
Troubling 825 0 0 0 0 0 0 0
ZSL = Zero-shot Learning, FSL = Few-shot Learning, CoT = Chain-of-Thought, ToT = Tree-of-Thought, GToT = Guided Tree-of-Thought.
Finding 12: FSL introduces the most critical bugs, while ZSL produces the safest but least functional test cases. GToT improves test structure but increases subtle defects, requiring additional validation.
4.4 [RQ4]: Readability of LLM-Generated Test Suites: Alignment with Developer Expectations
[Experiment Goal]: This experiment examines whether LLM-generated test suites meet developer readability expectations compared to EvoSuite. We focus on clarity, maintainability, and formatting. Using an automated readability model correlated with human evaluation, we assess how prompt engineering impacts readability compared to EvoSuite tests.
[Experiment Design]: We employ Scalabrino et al. (2018) readability model, trained on Java code including JUnit tests. This model evaluates syntactic structure, identifier clarity, textual coherence, and structural readability metrics, providing assessment beyond formatting-based analysis. We evaluate across three datasets (SF110, Defects4J, CMD) using multiple prompting techniques (ZSL, FSL, CoT, ToT, GToT), extracting readability scores for each test suite and aggregating results by model and strategy. Higher scores indicate more structured, easier-to-interpret tests—essential for real-world development. We also compare score distributions to identify approaches yielding improved structure and reduced complexity.
[Experiment Results]:
Dataset-Level Readability Analysis. Table 11 shows LLM-generated test suites consistently achieve higher readability than EvoSuite across all datasets. Readability varies with codebase complexity: Defects4J exhibits highest readability scores (GPT-3.5-Turbo: 81.73%, Mistral 7B: 69.06% vs. EvoSuite: 57.33%), followed by SF110 (GPT-3.5-Turbo: 75.95%, Mistral 7B: 70.28% vs. EvoSuite: 50.67%). CMD presents lowest scores (GPT-3.5-Turbo: 60.93%, Mistral 7B: 57.13%), indicating modern, evolving codebases challenge LLMs more. While LLMs consistently outperform EvoSuite in readability, their effectiveness varies with dataset complexity.
Finding 13: LLM-generated tests outperform EvoSuite in readability. However, readability varies with dataset complexity, as shown in Table 11. CMD, which contains modern and evolving codebases, presents the greatest challenge, yielding the lowest readability scores across all models.
Table 11: Statistical Metrics from Readability Score Distribution by Model and Dataset
Model Dataset Min Q1 Median Q3 Max Mean
GPT 3.5-turbo Defects4J 33.35 70.53 84.03 95.74 99.04 81.73
SF110 25.02 66.32 76.12 87.10 98.64 75.95
CMD 41.34 55.53 61.30 65.98 90.45 60.93
Mistral7B Defects4J 15.20 61.23 68.44 78.03 97.58 69.06
SF110 6.58 60.19 71.04 80.99 98.59 70.28
CMD 13.60 38.68 52.68 84.24 95.07 57.13
EvoSuite Defects4J 16.10 44.67 56.04 72.04 95.70 57.33
SF110 16.33 36.83 50.03 61.81 89.67 50.67
LLM vs. EvoSuite: Readability Performance. Table 12 shows LLM-generated tests consistently surpass EvoSuite in readability. For GPT-3.5-Turbo, GToT achieves the highest score (84.75% vs. EvoSuite’s 50.77%), followed by CoT (82.77%) and ToT (73.32%), while ZSL (77.93%) and FSL (73.49%) score lower. Similarly, for Mistral 7B, GToT leads (76.35%), followed by ToT (69.64%), with ZSL (63.79%) and FSL (60.40%) producing lowest scores. These results demonstrate reasoning-based prompting’s importance for readability.
Finding 14: LLM-generated test suites improve readability over EvoSuite by 21–40%, as shown in Table 12. This highlights the significant advantage of LLMs in producing more comprehensible test code.
Table 12: Statistical Metrics from Readability Score Distribution (Defects4J and SF110)
(a) GPT 3.5-turbo by Prompt vs EvoSuite
Prompt Min Q1 Median Q3 Max Mean
ZSL 29.37 66.79 77.50 93.50 98.41 77.93
FSL 25.02 60.24 74.36 88.07 97.78 73.49
CoT 43.28 73.17 85.74 95.33 99.04 82.77
ToT 33.36 64.33 72.04 82.64 97.83 73.32
GToT 42.24 75.20 89.39 96.26 98.64 84.75
EvoSuite 4.11 36.96 51.02 62.59 97.64 50.77
(b) Mistral7B by Prompt vs EvoSuite
Prompt Min Q1 Median Q3 Max Mean
ZSL 25.67 57.52 64.41 71.87 90.62 63.79
FSL 6.58 49.03 62.53 70.78 97.58 60.40
CoT 34.54 57.67 65.39 73.92 97.77 65.92
ToT 33.30 60.94 69.00 78.29 97.67 69.64
GToT 15.20 67.47 77.24 89.02 98.59 76.35
EvoSuite 4.11 36.96 51.02 62.59 97.64 50.77
Prompting Strategies: Influence on Readability. Table 13 shows reasoning-based prompting significantly enhances test suite clarity. GToT and CoT consistently achieve highest readability scores across datasets. For GPT-3.5-Turbo on CMD, CoT leads (67.99%), followed by GToT (61.86%), while ZSL (57.67%) and FSL (55.76%) score lower. Similarly, for GPT-4 on CMD, GToT achieves highest readability (67.71%). These results confirm minimal prompting provides insufficient guidance for well-structured tests, while reasoning-based approaches are essential for producing maintainable test suites.
Table 13: Statistical Metrics from Readability Score Distribution on CMD
(a) GPT 3.5-turbo on CMD
Prompt Min Q1 Median Q3 Max Mean
CoT 45.92 62.59 66.37 71.95 90.45 67.99
FSL 42.30 49.14 55.80 61.70 66.07 55.76
ToT 47.74 57.77 62.09 64.61 71.10 61.38
GToT 52.07 57.29 61.02 67.15 70.12 61.86
ZSL 41.34 53.19 57.95 63.26 69.17 57.67
(b) GPT 4 on CMD
Prompt Min Q1 Median Q3 Max Mean
CoT 51.51 57.83 61.42 66.12 75.20 61.64
FSL 30.14 47.15 50.56 56.03 71.80 51.62
ToT 54.02 59.94 62.86 66.51 73.04 63.18
GToT 58.10 63.35 68.85 71.08 76.13 67.71
ZSL 36.69 47.61 50.20 57.81 71.66 52.13
(c) Mistral7B on CMD
Prompt Min Q1 Median Q3 Max Mean
CoT
FSL 35.18 35.18 35.18 35.18 35.18 35.18
ToT 49.18 49.18 49.18 49.18 49.18 49.18
GToT 13.60 45.53 74.89 93.97 95.07 64.61
ZSL
(d) Mixtral8x7B on CMD
Prompt Min Q1 Median Q3 Max Mean
CoT 46.46 51.88 57.29 62.71 68.12 57.29
FSL 48.36 48.36 48.36 48.36 48.36 48.36
ToT 57.14 77.90 79.08 92.66 92.85 79.93
GToT 44.32 58.67 75.75 89.30 93.00 72.21
ZSL 38.43 44.17 49.91 55.65 61.39 49.91
  • 1
    Since only compilable code was considered, there was no compilable code for Mistral 7B’s CoT and ZSL prompts.
Finding 15: CoT and GToT consistently enhance readability, even on complex datasets like CMD, where unstructured prompting results in lower scores.
Model-Specific Readability Trends. Table 13 reveals how different LLMs respond to reasoning-based prompting. GPT-3.5-Turbo performs best with CoT (67.99%), while GPT-4 excels with GToT (67.71%), confirming larger models benefit from structured reasoning. Mistral 7B shows higher variability, reaching 64.61% with GToT, indicating sensitivity to prompt design. Mixtral 8x7B performs exceptionally with ToT (79.93%), though its limited sample size affects stability.
Finding 16: Model architecture, parameter size, and prompting strategy significantly impact readability, with larger models consistently improving when guided by reasoning-based prompting.
4.5 [RQ5]: Code Coverage.
[Experiment Goal]: This experiment evaluates LLM-generated test suite effectiveness in achieving structural coverage. We measure line, instruction, and method coverage, comparing LLMs against EvoSuite and examining how prompting strategies impact coverage. Method coverage is particularly important in our class-level evaluation to determine cross-method interaction rather than isolated functionality testing.
[Experiment Design]: We use JaCoCo [^12] https://www.jacoco.org/jacoco/ to measure three metrics: line coverage (executed source code lines), instruction coverage (executed bytecode instructions), and method coverage (methods invoked at least once). Following Tang et al. (2024), we exclude branch coverage as it only considers conditional branches and often yields missing values. We evaluate all compilable test cases across five prompting strategies (ZSL, FSL, CoT, ToT, GToT) against EvoSuite-generated tests across SF110, Defects4J, and CMD datasets. We compute statistical measures (min, max, quartiles, median, mean) to comprehensively analyze structural coverage.
[Experiment Results]:
EvoSuite vs. LLMs: Overall Coverage Performance. Tables 14, 15, and 16 reveal a significant performance gap between LLMs and EvoSuite. EvoSuite consistently outperforms LLMs across all metrics, achieving median coverage of 94.34% (line), 95.0% (instruction), and 100% (method) through search-based optimization. LLM-generated tests show higher variability, with reasoning-based prompting significantly improving coverage over ZSL and FSL. GPT-3.5-Turbo with CoT achieves the highest LLM median coverage (75.47% line, 76.0% instruction, 100% method), though all LLM configurations fall short of EvoSuite’s systematic exploration capabilities.
Table 14: Statistical metrics from Line coverage distribution.
(a) EvoSuite vs GPT 3.5-turbo.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 1.49 27.68 100.0 100.0 46.3
FSL 0.0 1.73 16.14 100.0 100.0 41.2
CoT 0.0 3.33 75.47 100.0 100.0 55.56
ToT 0.0 1.82 24.39 93.55 100.0 41.96
GToT 0.0 4.65 41.54 100.0 100.0 46.7
EvoSuite 0.0 88.61 94.34 97.67 100.0 91.75
(b) EvoSuite vs Mistral 7B.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 0.0 9.2 67.27 100.0 31.51
FSL 0.0 1.97 18.45 63.04 100.0 31.58
CoT 0.0 0.0 1.12 65.22 100.0 30.28
ToT 0.0 1.2 20.0 76.44 100.0 37.0
GToT 0.0 4.55 30.0 83.89 100.0 42.41
EvoSuite 0.0 88.61 94.34 97.67 100.0 91.75
(c) GPT 3.5-turbo by datasets.
Dataset Min Q1 Median Q3 Max Mean
Defects4J 0.0 10.22 82.74 100.0 100.0 62.61
SF110 0.0 0.0 3.57 23.24 100.0 22.72
CMD 4.0 53.59 100.0 100.0 100.0 76.86
(d) Mistral 7B by datasets.
Dataset Min Q1 Median Q3 Max Mean
Defects4J 0.0 4.38 63.4 89.47 100.0 52.55
SF110 0.0 0.0 3.03 25.0 100.0 15.15
CMD 0.0 0.0 0.0 0.0 0.0 0.0
(e) GPT 3.5-turbo on CMD.
Prompt Min Q1 Median Q3 Max Mean
ZSL 4.0 22.88 61.29 92.31 100.0 56.05
FSL 8.33 30.0 83.33 100.0 100.0 67.78
CoT 8.33 92.05 100.0 100.0 100.0 84.53
ToT 31.82 73.39 100.0 100.0 100.0 85.84
GToT 31.82 67.8 100.0 100.0 100.0 84.27
(f) GPT 4 on CMD.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 0.0 8.33 70.97 89.47 31.88
FSL 5.56 12.89 52.51 85.94 100.0 52.74
CoT 0.0 18.68 65.22 100.0 100.0 60.35
ToT 0.0 28.25 72.22 100.0 100.0 63.82
GToT 0.0 31.61 72.22 100.0 100.0 64.01
Finding 17: EvoSuite outperforms all LLMs in coverage metrics. However, reasoning-based techniques (CoT, GToT, ToT) significantly enhance LLM test effectiveness, reducing the performance gap.
Dataset-Level Coverage Trends. Tables 1415, and 16 show coverage varies significantly across datasets. Defects4J achieves highest coverage (GPT-3.5-Turbo: 82.74% line, 85.0% instruction, 100% method), suggesting its structured design facilitates test generation with clearer functional boundaries. In contrast, SF110 reports much lower coverage (3.57% line, 2.0% instruction, 16.67% method) due to higher complexity and diverse architectures. CMD reports highest median coverage (100% for some prompts), though this likely reflects its smaller size and modular structure rather than superior LLM capabilities. The contrast between datasets highlights structural influences: Defects4J benefits from potential training exposure (Sallou et al. (2023)), clearer functional boundaries, and modularized code, while SF110 presents challenges through higher complexity and lower public accessibility. Finding 18: LLMs achieve higher coverage on well-documented datasets like Defects4J but struggle with complex projects like SF110, where greater structural variability makes test generation more challenging. CMD’s high coverage is likely due to its small, modular dataset rather than an inherent LLM advantage.
Table 15: Statistical metrics from Instruction coverage distribution.
(a) EvoSuite vs GPT 3.5-turbo.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 1.0 29.0 100.0 100.0 46.26
FSL 0.0 1.0 14.0 100.0 100.0 41.72
CoT 0.0 3.0 76.0 100.0 100.0 55.64
ToT 0.0 1.0 27.0 96.0 100.0 42.91
GToT 0.0 3.5 45.0 100.0 100.0 47.61
EvoSuite 0.0 91.0 95.0 98.0 100.0 93.35
(b) EvoSuite vs Mistral 7B.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 0.0 10.0 73.25 100.0 33.63
FSL 0.0 1.0 18.0 60.5 100.0 32.48
CoT 0.0 0.0 0.0 83.0 100.0 31.19
ToT 0.0 0.0 17.0 81.0 100.0 37.49
GToT 0.0 3.25 26.0 87.25 100.0 42.31
EvoSuite 0.0 91.0 95.0 98.0 100.0 93.35
(c) GPT 3.5-turbo by datasets.
Dataset Min Q1 Median Q3 Max Mean
Defects4J 0.0 11.0 85.0 100.0 100.0 63.14
SF110 0.0 0.0 2.0 24.5 100.0 22.58
CMD 2.0 48.25 100.0 100.0 100.0 76.56
(d) Mistral 7B by datasets.
Dataset Min Q1 Median Q3 Max Mean
Defects4J 0.0 2.25 66.0 93.0 100.0 53.62
SF110 0.0 0.0 2.0 22.0 100.0 15.37
CMD 0.0 0.0 0.0 0.0 0.0 0.0
(e) GPT 3.5-turbo on CMD.
Prompt Min Q1 Median Q3 Max Mean
ZSL 2.0 22.5 60.0 91.0 100.0 55.27
FSL 9.0 31.0 81.0 100.0 100.0 67.18
CoT 7.0 93.75 100.0 100.0 100.0 84.65
ToT 28.0 77.25 100.0 100.0 100.0 85.65
GToT 28.0 69.75 100.0 100.0 100.0 84.06
(f) GPT 4 on CMD.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 0.0 6.0 61.0 95.0 31.46
FSL 3.0 15.5 51.5 82.25 100.0 51.92
CoT 0.0 18.0 61.0 100.0 100.0 61.05
ToT 0.0 29.0 79.5 100.0 100.0 65.19
GToT 0.0 29.5 79.5 100.0 100.0 65.21
While Sallou et al. (2023) attributes poor SF110 performance mainly to data accessibility, our analysis demonstrates SF110’s inherent complexity complicates test generation (see Table 2 in Section 3.6). However, Defects4J’s broad accessibility raises data leakage concerns, as LLMs may have encountered its test cases during pretraining. As shown in (Sallou et al., 2023), ChatGPT can retrieve detailed information about specific Defects4J bugs, suggesting performance may reflect memorization rather than genuine reasoning.
Finding 19: Dataset characteristics influence LLM performance—Defects4J’s high coverage likely stems from pretraining exposure while SF110’s limited representation restricts generalization, highlighting the need for careful dataset selection to mitigate data leakage in test generation studies.
Impact of Reasoning-Based Prompting on Coverage. Tables 141516 show reasoning-based techniques (CoT, GToT, ToT) consistently improve coverage compared to ZSL and FSL. For line coverage, GPT-3.5-Turbo with CoT achieves highest median (75.47%), while ZSL and FSL produce significantly lower coverage (27.68% and 16.14%). Mistral 7B performs best with GToT (30.0%), though still below GPT-3.5-Turbo. For instruction coverage, EvoSuite maintains advantage (95.0%). Among LLMs, GPT-3.5-Turbo with CoT (76.0%) and GToT (45.0%) lead, while Mistral 7B’s best result with GToT (26.0%) indicates smaller models’ limitations. For method coverage, GPT-3.5-Turbo with CoT reaches 100% in some cases, while Mistral 7B improves with GToT (89.59%) but still lags. These results demonstrate reasoning-based prompting enhances LLM performance significantly but doesn’t fully bridge the gap with EvoSuite.
Table 16: Statistical metrics from Method coverage distribution.
(a) EvoSuite vs GPT 3.5-turbo.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 6.25 83.33 100.0 100.0 56.95
FSL 0.0 7.14 66.67 100.0 100.0 55.31
CoT 0.0 14.29 100.0 100.0 100.0 65.18
ToT 0.0 8.33 66.67 100.0 100.0 55.31
GToT 0.0 16.67 76.92 100.0 100.0 60.32
EvoSuite 0.0 100.0 100.0 100.0 100.0 99.21
(b) EvoSuite vs Mistral 7B.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 0.0 29.16 100.0 100.0 48.83
FSL 0.0 9.94 72.22 100.0 100.0 55.1
CoT 0.0 0.0 6.9 97.06 100.0 41.31
ToT 0.0 5.96 50.0 100.0 100.0 55.07
GToT 0.0 17.5 89.59 100.0 100.0 62.56
EvoSuite 0.0 100.0 100.0 100.0 100.0 99.21
(c) GPT 3.5-turbo by datasets.
Dataset Min Q1 Median Q3 Max Mean
Defects4J 0.0 33.33 100.0 100.0 100.0 73.86
SF110 0.0 0.0 16.67 89.44 100.0 35.17
CMD 20.0 100.0 100.0 100.0 100.0 89.35
(d) Mistral 7B by datasets.
Dataset Min Q1 Median Q3 Max Mean
Defects4J 0.0 17.05 100.0 100.0 100.0 70.1
SF110 0.0 0.0 20.0 50.0 100.0 34.7
CMD 0.0 0.0 0.0 0.0 0.0 0.0
(e) GPT 3.5-turbo on CMD.
Prompt Min Q1 Median Q3 Max Mean
ZSL 20.0 35.41 90.91 100.0 100.0 71.29
FSL 25.0 88.89 100.0 100.0 100.0 84.15
CoT 20.0 100.0 100.0 100.0 100.0 91.0
ToT 75.0 100.0 100.0 100.0 100.0 98.04
GToT 75.0 100.0 100.0 100.0 100.0 97.82
(f) GPT 4 on CMD.
Prompt Min Q1 Median Q3 Max Mean
ZSL 0.0 0.0 40.0 75.0 100.0 43.56
FSL 16.67 38.33 65.0 100.0 100.0 64.86
CoT 0.0 34.72 100.0 100.0 100.0 69.8
ToT 0.0 62.5 100.0 100.0 100.0 79.59
GToT 0.0 70.46 100.0 100.0 100.0 80.25
Finding 20: Reasoning-based prompting significantly enhances LLM coverage across all datasets and models. However, smaller models like Mistral 7B exhibit greater limitations even with reasoning-based prompting.
4.6 [RQ6]: Execution Outcomes and Fault-Detection Capabilities
[Experiment Goal]: While coverage indicates broad execution, it doesn’t guarantee usefulness or fault detection. This experiment evaluates the execution outcomes and fault detection potential, to determine whether LLM-generated tests can effectively contribute to quality assurance
[Experiment Design]: We analyze execution behavior across three datasets (SF110, Defects4J, CMD) and multiple prompting techniques. Tests are classified as Passed (executes successfully), Failed (compilable but fails at runtime), or Timeout (exceeds 4-second limit, matching EvoSuite). We conduct failure analysis, categorizing runtime errors (ClassNotFoundException, IllegalArgumentException, NullPointerException, etc.) to distinguish flawed test logic from genuine fault detection. To assess fault-detection capabilities, we use Pitest [^13] https://github.com/hcoles/pitest to introduce controlled faults, measuring the mutation score (percentage of killed mutants) as a quantitative measure of defect identification effectiveness.
[Experiment Results]:
Test Execution Outcomes: Pass, Fail, and Timeout Rates. Our analysis reveals significant performance variations across test generation approaches. Table 17(a) shows GPT-3.5-Turbo with notable failure rates (77.11% for Defects4J, 86.75% for SF110) while Mistral 7B and Mixtral 8x7B fail 100% of CMD tests. Table 17(b) demonstrates Chain-of-Thought prompting reduces GPT-3.5-Turbo’s failures to 64.40% on Defects4J, showing clear benefits over Zero-Shot and Few-Shot approaches. Table 17(c) reveals similar CoT benefits on SF110, while Table 17(d) highlights industrial codebases’ challenges on CMD. Timeout rates reflect test complexity differences, with EvoSuite showing the highest rates (34.47% in SF110) due to its complex execution paths, while LLM-generated tests have significantly lower timeout rates (GPT-3.5-Turbo: 6.95% in SF110, 8.22% in Defects4J). Pass rates also reveal trade-offs, with GPT-3.5-Turbo achieving moderate rates (6.31% in SF110, 14.68% in Defects4J) compared to EvoSuite’s intentionally lower rates (0.27% in SF110, 2.51% in Defects4J) which align with its fault-revealing objectives. These execution metrics reveal a crucial challenge in test generation evaluation: without a reliable test oracle, distinguishing between genuine fault detection and flawed test logic remains difficult, while higher pass rates or lower timeout rates don’t necessarily indicate superior test effectiveness.
Table 17: Summary of test execution statuses (passed, failed, timed out) for LLM-generated test suites and EvoSuite.
(a) By LLM and EvoSuite across all datasets.
Models/Tool SF110 Defects4J CMD
FAILED PASSED TIMEOUT FAILED PASSED TIMEOUT FAILED PASSED TIMEOUT
GPT 3.5-turbo 86.75% 6.31% 6.95% 77.11% 14.68% 8.22% 61.66% 36.15% 2.19%
GPT 4 82.65% 5.99% 11.36%
Mistral7B 76.57% 0.85% 22.59% 86.05% 7.16% 6.80% 100% 0% 0%
Mixtral8x7B 100% 0% 0%
EvoSuite 62.26% 0.27% 34.47% 91% 2.51% 6.49%
(b) GPT 3.5-turbo and Mistral7B on Defects4J, by Prompt.
Models FAILED (%) PASSED (%) TIMEOUT (%)
ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2
GPT 3.5-turbo 75.75 75.64 64.40 84.65 85.09 14.40 13.91 29.25 5.70 10.40 10.12 10.45 6.36 9.65 4.51
Mistral7B 83.53 82.98 90.79 87.16 85.77 10.59 16.31 1.32 3.38 4.18 5.88 0.71 7.89 9.46 10.04
(c) GPT 3.5-turbo and Mistral7B on SF110, by Prompt.
Model Name FAILED (%) PASSED (%) TIMEOUT (%)
ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2
GPT 3.5-turbo 86.63 89.55 78.84 89.70 89.01 7.67 3.52 14.56 1.94 3.85 5.71 6.92 6.60 8.36 7.14
Mistral7B 72.67 97.39 41.51 83.89 87.38 0 0 0 2.78 1.46 27.33 2.61 58.49 13.33 11.17
(d) All LLMs on CMD, by Prompt.
Models FAILED (%) PASSED (%) TIMEOUT (%)
ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2
GPT 3.5-turbo 96.49 65.05 33.60 52.21 60.96 2.63 29.13 64.40 46.90 37.67 0.88 5.83 2.00 0.88 1.37
GPT 4 68.12 94.19 80.58 86.49 83.87 0 1.16 7.77 8.11 12.90 31.88 4.65 11.65 5.41 3.23
Mistral7B 100 100 100 0 0 0 0 0 0
Mixtral8x7B 100 100 100 100 0 0 0 0 0 0 0 0
Finding 21: Test execution metrics reveal complex trade-offs in automated test generation—high failure rates (33.60-100%) reflect challenges in distinguishing genuine faults from flawed tests, lower timeout rates may indicate simpler rather than better tests, and higher pass rates don’t necessarily demonstrate superior fault detection capability.
Failure Analysis: Error Types and Root Causes. Our analysis (Table 18 and Figure LABEL:fig:failling-test-heatmap) reveals that most LLM-generated test suites fail during execution, with three predominant error types, with the y-axis of Figure LABEL:fig:failling-test-heatmap dynamically displaying error types unique to each model and test suite configuration. ClassNotFoundException (CNFE) dominates failures in larger models (GPT-3.5-Turbo: 91.76% in Defects4J, 93.89% in SF110), while IllegalArgumentException (IAE) significantly affects smaller models like Mistral 7B (22.02% in Defects4J) and both Mistral models on CMD (50%). Manual inspection confirmed these failures stem from hallucinated dependencies and invalid API usage rather than successful bug detection. Prompting strategies create notable trade-offs—ZSL and FSL generate the highest CNFE rates (GPT-4/CMD: 98.0% with ZSL vs. 75.0% with GToT), while reasoning-based approaches reduce hallucinations but increase AssertionErrors and NullPointerExceptions. Model size directly affects error patterns, with smaller models showing context-dependent failure distributions that highlight their limited repository understanding.
Table 18: Errors from failing tests by model and dataset
Model Dataset AE CNFE FNFE IOE IAE ISE IOOBE MUE NPE NFE OOME RE
GPT 3.5-Turbo Defects4J 0.88 91.76 0.02 2.91 2.92 0.47 0.56 0.02 0.38 0.03 0 0.04
SF110 0.43 93.89 0.20 0.04 4.25 0.48 0.06 0 0.48 0.07 0.03 0.04
CMD 2.60 94.80 0 0 1.20 0 0 0 1.20 0.20 0 0.00
GPT 4 CMD 5.83 83.18 0 0 2.99 0 0 0 5.70 2.16 0 0.15
Mistral 7B Defects4J 9.59 54.93 0.00 2.62 22.02 2.80 1.99 0.21 5.49 0.34 0 0.00
SF110 3.77 59.98 0.00 0.00 19.13 7.98 0.57 0 7.46 0.93 0.00 0.17
CMD 0.00 50.00 0 0 50.00 0 0 0 0.00 0.00 0 0.00
Mixtral 8x7B CMD 0.00 50.00 0 0 50.00 0 0 0 0.00 0.00 0 0.00
  • 1
    AE: AssertionError, CNFE: ClassNotFoundException, FNFE: FileNotFoundException, IOE: IOError, IAE: IllegalArgumentException, ISE: IllegalStateException, IOOBE: IndexOutOfBoundsException, MUE: MalformedURLException, NPE: NullPointerException, NFE: NumberFormatException, OOME: OutOfMemoryError, RE: RuntimeException.
Finding 22: LLM test generation failures stem primarily from model limitations rather than successful bug detection. Three key error patterns dominate: hallucinated dependencies (CNFE, 83-94% in larger models), incorrect argument validation (IAE, 19-50% in smaller models), and assertion mismatches (higher in reasoning-based models). While reasoning-based prompting reduces hallucinations, it increases assertion errors, revealing a fundamental trade-off between dependency resolution and assertion quality that varies with model architecture.
Mutation testing. Many ”Passed” test cases targeted abstract classes, interfaces, or method skeletons lacking executable logic, resulting in 0% line coverage and 0% mutation coverage in Pitest. While these tests did not produce errors, they failed to execute meaningful code, making them ineffective for fault detection. Coverage analysis in Section 4.5 further confirms this issue, with median line coverage as low as 3.57% (GPT-3.5-Turbo on SF110) and 3.03% (Mistral 7B on SF110) (Table 14). Given that most test suites were either empty (only containing test class definitions) or targeting non-executable components, computing mutation scores per class or method would not provide relevant insights.
Finding 23: Mutation testing was largely ineffective due to the non-executable nature of LLM-generated passed tests, which often targeted abstract classes or contained empty test definitions. Code coverage results confirm their failure to interact meaningfully with the code.
4.7 [RQ7]: Test Smell detection.
[Experiment Goal]: Our aim is to evaluate the extent to which the test suites generated by LLMs exhibit test smells compared with those produced by an SBST method like EvoSuite.
[Experiment Design]: This experiment assesses test smell prevalence using TsDetect (Peruma et al., 2020), which identifies 20 different smell types (Peruma et al., 2019, 2020). We record the proportion of test suites containing at least one instance of each smell type. These percentages reflect smell distribution rather than absolute counts, enabling comparative analysis of structural quality and maintainability.
[Experiment Results]:
Test Smell Distribution Across LLMs and EvoSuite. Table 19 reveals Assertion Roulette (AR), Conditional Logic Test (CLT), and Magic Number Test (MNT) as the most prevalent smells. MNT appears in nearly 100% of test suites across models, confirming Siddiq et al. (2024a)’s observation of MNT’s ubiquity. Our AR variability (73.15% with GPT-3.5/FSL on Defects4J, dropping to 19.86% with GToT on CMD) aligns with their reported range (4.6–61.3%). EH is more prevalent in EvoSuite (93.10%) than in LLMs (10–30%), partly contradicting Siddiq et al. (2024a)’s reported absence of EH in LLM outputs. This suggests reasoning-aware prompts may trigger more robust exception modeling, consistent with Finding 22. We report absence of Default Test, Verbose Test, and Dependent Test smells, supporting Finding 21 that LLM tests tend to be structurally simpler. The increase in Empty Test (EM) rates with reasoning prompts complements Finding 17, showing structured prompting improves formatting but may lead to shallow content. While Siddiq et al. (2024a) briefly report EM smells, we expand by demonstrating how prompt complexity influences EM rates across datasets.
Table 19: Test smells distribution by Model, Prompt and by Datasets.
Model Dataset Prompt Test Smell (%)
AR CLT CI EM EH GF MG RP RA SE ST EA LT DA UT IT RO MNT
GPT 3 SF110 ZSL 46.17 2.68 0.16 12.71 13.02 11.13 2.92 0.0 0.79 23.52 0.39 48.38 53.75 9.31 27.70 0.0 3.24 99.84
FSL 74.44 3.91 0.11 7.14 14.51 6.70 3.57 0.0 3.13 33.15 0.45 70.65 53.91 16.52 16.29 0.0 4.13 100
CoT 40.98 2.15 0.0 22.25 11.60 5.93 1.98 0.0 0.95 17.53 0.09 42.44 50.26 6.53 34.54 0.0 1.89 99.05
ToT 59.77 3.22 0.11 4.94 19.20 13.45 2.41 0.34 3.56 28.51 0.92 56.90 76.21 13.33 15.06 0.11 2.64 100
GToT 38.23 0.58 0.0 28.97 10.24 2.77 1.99 0.0 2.13 11.11 0.05 42.84 40.03 3.30 37.17 0.0 2.23 100
Defects4J ZSL 57.01 4.05 0.06 15.72 24.30 7.50 1.09 0.0 0.91 26.90 0.79 48.55 60.70 9.92 19.65 0.0 1.09 99.64
FSL 73.15 4.90 0.24 13.18 24.68 3.94 0.88 0.0 2.09 31.19 1.61 64.95 57.64 17.36 15.19 0.0 0.96 100
CoT 40.06 2.83 0.11 34.40 16.79 2.99 0.61 0.0 0.66 17.73 0.61 36.45 46.81 5.71 41.72 0.0 0.66 99.28
ToT 69.56 5.52 0.32 4.57 30.60 11.20 1.03 0.0 0.79 29.81 0.79 52.92 75.95 11.20 6.23 0.0 1.0 99.97
GToT 29.64 2.06 0.0 52.39 12.67 2.03 0.90 0.0 1.11 7.97 0.24 28.42 32.22 1.45 55.05 0.0 1.0 99.97
CMD ZSL 37.72 0.0 0.0 33.33 15.79 35.96 0.0 0.0 0.0 2.63 0.0 30.70 27.19 0.0 54.39 0.0 0.0 100
FSL 18.45 0.97 0.0 49.51 25.24 14.56 0.0 0.0 0.0 0.0 0.0 18.45 16.50 0.0 77.67 0.0 0.0 100
CoT 6.4 0.0 0.0 78.80 7.20 2.80 0.0 0.0 0.40 0.0 0.0 6.40 7.20 0.0 89.20 0.0 0.0 100
ToT 15.04 0.0 0.0 71.68 7.96 18.58 0.0 0.0 0.0 0.88 0.0 17.70 15.04 0.88 76.99 0.0 0.0 100
GToT 19.86 0.0 0.0 58.90 15.07 17.81 0.0 0.0 0.68 1.37 0.0 18.49 21.23 0.0 74.66 0.0 0.0 100
GPT 4 CMD ZSL 30.43 1.45 0.0 11.59 14.49 20.29 0.0 0.0 1.45 2.90 0.0 36.23 36.23 0.0 34.78 0.0 0.0 98.55
FSL 41.18 1.18 0.0 10.59 31.76 20.0 0.0 1.18 0.0 2.35 0.0 48.24 45.88 2.35 36.47 0.0 0.0 100
CoT 14.56 0.0 0.0 28.16 18.45 17.48 0.0 0.0 0.0 0.97 0.0 18.45 35.92 0.0 54.37 0.0 0.0 98.06
ToT 16.90 0.0 0.0 33.80 16.90 15.49 0.0 0.0 0.0 0.0 0.0 23.94 22.54 0.0 53.52 0.0 0.0 97.18
GToT 13.98 0.0 0.0 49.46 17.20 19.35 0.0 0.0 1.08 0.0 0.0 16.13 21.51 0.0 69.89 1.08 0.0 97.85
Mistral 7B SF110 ZSL 63.45 9.66 0.0 0.0 14.71 7.98 0.42 1.26 1.26 31.51 0.84 65.13 80.25 10.08 3.36 0.42 0.42 100
FSL 72.69 11.89 0.0 0.88 7.05 0.44 0.0 1.32 6.17 27.75 0.0 68.28 69.60 15.86 5.29 0.0 0.0 100
CoT 52.26 7.04 0.0 2.51 19.10 2.01 0.5 1.51 1.51 22.61 0.5 53.27 84.92 11.56 5.53 0.0 0.5 94.47
ToT 42.59 2.65 0.0 8.47 12.70 2.38 0.26 0.26 1.59 10.58 0.0 43.12 42.59 8.99 11.11 0.0 0.26 92.59
GToT 21.20 0.82 0.14 13.54 6.57 0.82 0.14 0.0 0.55 3.42 0.0 27.91 16.01 2.74 21.34 0.0 0.14 85.64
Defects4J ZSL 73.64 14.34 0.39 0.0 17.44 1.55 0.78 0.78 1.55 13.18 1.55 48.06 58.14 32.17 2.33 0.0 0.39 100
FSL 79.35 14.17 0.4 1.62 8.5 0.81 0.4 0.0 1.21 25.51 0.81 62.75 78.95 21.46 2.02 0.4 0.4 100
CoT 65.13 13.82 0.0 0.66 21.05 0.66 0.0 0.0 1.32 10.53 1.32 41.45 81.58 31.58 0.66 0.0 0.0 95.39
ToT 69.84 7.54 0.4 1.59 22.62 2.38 0.0 0.0 1.59 22.22 1.59 64.29 78.97 23.41 3.17 0.0 0.0 98.41
GToT 27.97 2.89 0.0 12.86 13.67 0.0 0.0 0.0 0.16 4.82 0.16 28.94 30.55 4.18 16.08 0.0 0.0 84.41
CMD ZSL - - - - - - - - - - - - - - - - - -
FSL 0.0 0.0 0.0 0.0 0.0 100 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 100 0.0 0.0 100
CoT - - - - - - - - - - - - - - - - - -
ToT 0.0 0.0 0.0 0.0 100 0.0 0.0 0.0 0.0 0.0 0.0 0.0 100 0.0 100 0.0 0.0 100
GToT 16.67 0.0 0.0 33.33 0.0 0.0 0.0 0.0 0.0 0.0 0.0 16.67 0.0 0.0 33.33 0.0 0.0 66.67
Mixtral8x7B CMD ZSL 75 0.0 0.0 0.0 25.0 75.0 0.0 0.0 0.0 0.0 0.0 25.0 25.0 0.0 0.0 0.0 0.0 100
FSL 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 33.33 0.0 0.0 33.33 0.0 0.0 66.67
CoT 0.0 0.0 0.0 0.0 50 0.0 0.0 0.0 0.0 0.0 0.0 0.0 50.0 0.0 50.0 0.0 0.0 100
ToT 44.44 11.11 0.0 44.44 11.11 0.0 0.0 0.0 0.0 11.11 0.0 55.56 33.33 0.0 55.56 0.0 0.0 100
GToT 0.0 0.0 0.0 50.0 50.0 50.0 0.0 0.0 0.0 0.0 0.0 0.0 50.0 0.0 100 0.0 0.0 100
EvoSuite SF110 50.78 0.0 0.0 6.90 93.10 0.0 4.44 0.0 0.0 17.53 0.0 54.67 76.94 0.07 6.90 0.0 3.96 100
Defects4J 38.24 0.0 0.0 6.90 93.10 0.0 4.44 0.0 0.0 10.71 0.0 62.95 82.18 6.73 0.38 0.0 4.34 100
  • 1
    AR: Assertion Roulette, CLT: Conditional Logic Test, CI: Constructor Initialization, EM: Empty Test, EH: Exception Handling, GF: General Fixture, MG: Mistery Guest, RP: Redundant Print, RA: Redundant Assertion, SE: Sensitive Equality, ST: Sleepy Test, EA: Eager Test, LT: Lazy Test, DA: Duplicate Assert, UT: Unknown Test, IT: Ignored Test, RO: Resource Optimism, MNT: Magic Number Test, ZSL: Zero-shot Learning, FSL: Few-shot Learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
Finding 24: LLM-generated tests show simpler structures, with fewer complexity-related smells such as GF (0.0–35.96%), RA (0.0–6.17%), and CLT (0.0–14.34%). In contrast, EvoSuite consistently produces higher test smells, notably EH (93.10%), LT (76.94–82.18%), and EA (54.67–62.95%), reflecting its aggressive search-based optimization.
Dataset-Level Trends in Test Smell Prevalence. Test smell distribution reveals dataset-specific trends. Defects4J exhibits the highest Assertion Roulette rates (73.15% GPT-3.5-Turbo/FSL), indicating challenges in generating precise assertions for real-world bugs. Our findings partially align with Siddiq et al. (2024a)’s observations, but expand the analysis by comparing SF110, Defects4J, and CMD datasets. SF110 shows elevated General Fixture (93.10% in EvoSuite) and 100% Magic Number Test, while CMD demonstrates the lowest Assertion Roulette (6.4% GPT-3.5-Turbo/CoT). The analysis reveals that LLM-generated tests tend to be stateless and isolated, lacking the inter-method or state-dependent sequences typical of EvoSuite tests. This comprehensive examination illuminates how project context and software evolution influence test smell profiles across different datasets.
Finding 25: LLM-generated CMD tests show limited capturing of shared object states (GF: 0.0–35.96%, MG: 0.0%) compared to EvoSuite (4.44%), revealing challenges in modeling stateful interactions. Defects4J tests display high AR rates (27.97–79.35%), especially in FSL, underlining difficulties in generating precise assertions for for real-world defects localization.
Impact of Prompt Engineering on Test Quality. Test smell prevalence varies across LLMs, with CoT and GToT prompts improving quality. GPT-4 (CMD) reduces Assertion Roulette to 14.56%, and GPT-3.5-Turbo achieves the lowest AR at 6.4% (CoT on CMD), yet both still show nearly 100% Magic Number Test. Mistral 7B and Mixtral 8x7B yield more smells, with Mistral 7B reaching 27.97% AR on Defects4J. EvoSuite, though competitive, exhibits 100% MNT and notable AR rates (38.24–50.78%). These findings align with Siddiq et al. (2024a)’s report of universally high MNT and variable AR. Our results extend their analysis by evaluating more models and prompts, demonstrating how reasoning-based prompting reduces AR. This prompt-model-smell interplay emphasizes that improvements depend on structured prompting strategies, not just model capabilities. MNT’s persistence confirms its status as a core limitation in current test generation methods.
Finding 26: CoT and GToT produce the cleanest tests, notably reducing AR smells (GPT-3 CoT: 6.4–40.98%, GPT-3 GToT: 19.86–38.23%, GPT-4 CoT: 14.56%) compared to FSL (18.45–79.35%) and EvoSuite (38.24–50.78%). However, all strategies consistently struggle with Magic Number smells (85.64–100%), highlighting ongoing challenges with numeric literals.
5 Threats to Validity
This section addresses the threats to the validity of this study.
External Validity. The generalizability of our findings faces several limitations. Though we used three datasets (Defects4J, SF110, and CMD), these may not fully represent real-world software complexity. CMD’s small size of just two projects restricts our observations, particularly regarding the unexpected cases where GPT-3.5-Turbo outperformed GPT-4, suggesting model performance varies significantly with dataset characteristics and prompting approaches. Further experimentation on a broader dataset is necessary to confirm our observations.
Despite our efforts to mitigate training data contamination through CMD, LLMs may still have encountered similar test patterns during pretraining, potentially overestimating their true capabilities—a persistent challenge in LLM-based software engineering research. Our evaluation of four models from OpenAI and MistralAI further limits the breadth of our conclusions. Future work should explore a wider range of models, including open-source LLMs trained specifically for software engineering tasks.
Model knowledge cutoff dates also likely influenced our results, with Mistral 7B’s older knowledge base (2021-2022) compared to Mixtral 8x7B’s more recent release (December 2023) creating performance disparities. Our findings reflect these models’ capabilities during our study period (November 2023–April 2024), and newer versions may demonstrate different results.
Internal Validity. Several internal threats could affect our results’ reliability. LLM output randomness introduces variability in generated tests, which we mitigated by executing each prompt 30 times per test suite and performing statistical analyses. Similarly, we ran 30 iterations to address EvoSuite’s variability from its evolutionary search process.
Test smell detection methodology presents another challenge. Recent studies (Panichella et al., 2022) highlight concerns about detection tools’ reliability, noting high false-positive and false-negative rates. Some smells like Assertion Roulette and Eager Test don’t always indicate poor maintainability in LLM-generated tests. To mitigate this, we manually validated a subset of test cases to better assess whether detected smells genuinely impact maintainability, though this doesn’t eliminate all possible biases.
Our mutation testing analysis using Pitest faced limitations as only ”Passed” tests were analyzed. Most Passed tests targeted interfaces, abstract classes, or contained empty methods, leading to 0% line coverage and mutation score. This exposes the ineffectiveness of Passed LLM-generated tests for fault detection rather than indicating a failure of mutation testing methodology. The small dataset of Passed tests further limited our findings’ generalizability.
Finally, prompting strategy effectiveness represents another concern. While GToT consistently improves test structure and reliability, prompt designs may need further refinement for different models and datasets. Future research should explore model-specific prompt tuning to optimize LLM-generated test quality.
6 Related work
6.1 LLM-Based Test Generation: Strengths and Limitations
Recent research has explored the potential of Large Language Models (LLMs) for automated unit test generation, primarily focusing on correctness and code coverage. Siddiq et al. (2024a) examined ChatGPT-generated tests, emphasizing compilability and correctness issues, but their study did not evaluate key aspects such as maintainability, readability, or long-term usability. Similarly, Tang et al. (2024) and Yang et al. (2024) investigated coverage in LLM-generated tests, concluding that while LLMs can generate syntactically correct tests, they often fail at execution due to hallucinated dependencies and lack fault-detection capabilities.
Prior work focuses on static correctness, while we provide a comprehensive evaluation of LLM-generated tests. Our study assesses readability, maintainability, runtime failures, fault detection via mutation testing, and test smells, setting new benchmarks in software testing research.
6.2 Prompt Engineering for Test Generation
The impact of prompt engineering on LLM-based test generation has been explored in recent research. Yuan et al. (2024) and Chen et al. (2024) examined the effects of different prompting techniques on test correctness, demonstrating that Few-Shot Learning (FSL) improves correctness but does not fully mitigate hallucinated dependencies. Similarly, Yang et al. (2024) introduced Chain-of-Thought (CoT) prompting, highlighting its ability to enhance logical reasoning in test generation. However, these studies primarily focused on static correctness evaluations and did not analyze runtime failures, maintainability, or test smells, limiting their applicability in real-world testing workflows.
We systematically compare five prompting strategies—ZSL, FSL, CoT, ToT, and GToT—by executing generated tests and analyzing runtime outcomes (Pass, Fail, Timeout). Our study evaluates their impact on fault detection, test maintainability, and practical reliability, bridging the gap between theoretical correctness and real-world testing.
6.3 Execution-Based Correctness and Fault Detection
Prior works evaluating LLM-generated tests, such as Schafer et al.(Schäfer et al., 2023), Siddiq et al.(Siddiq et al., 2024a), and Chen et al. (Chen et al., 2024), have predominantly focused on static correctness criteria, including compilability and structural coverage. However, Palomba et al. (2016b) emphasized that high coverage does not necessarily imply meaningful tests—a concern validated by Bhatia et al. (2024), who found that while LLM-generated tests often execute successfully, they frequently lack effective fault-detection capabilities. To address this, we perform execution-based analysis by classifying test outcomes into Passed, Failed, and Timeout categories. We also apply mutation testing (Pitest) to evaluate fault detection—a previously overlooked aspect. Our findings show many LLM-generated tests include shallow assertions, limiting their defect detection effectiveness.
By going beyond static correctness checks, our study provides a real-world evaluation of test suite effectiveness, offering insights into critical weaknesses in LLM-generated test logic and execution reliability.
6.4 Readability, Maintainability, and Test Smell Analysis
While several prior studies acknowledge readability or maintainability issues in LLM-generated tests, they remain limited in scope. Siddiq et al. (2024a) mention readability concerns without structured evaluation, Yang et al. (2024) focus primarily on correctness metrics omitting test smell analysis, and Tang et al. (2024) and Yuan et al. (2023) report code style metrics without investigating maintainability impacts across diverse datasets.
Our study presents the first comprehensive multi-dimensional evaluation of LLM-generated test maintainability. We systematically analyze test smells using TsDetect, evaluate readability with a deep learning model tailored to JUnit code (Scalabrino et al., 2018), and assess structural conformance (MSR) and extraction feasibility (CSR)—dimensions not addressed in previous evaluations.
By integrating these metrics, our work uncovers maintainability challenges and provides actionable insights for prompt design and post-processing. This holistic evaluation addresses structural usability, setting a new benchmark for evaluating the practical viability of LLM-generated test suites.
6.5 Advancing LLM-Based Unit Test Improvement
Recent studies have explored LLM-driven unit test improvement techniques, primarily focusing on iterative refinement. Gu et al. (2024) introduced TestArt, a co-evolutionary approach that iteratively refines LLM-generated tests through repair iterations, enhancing correctness over multiple steps. Similarly, Alshahwan et al. (2024) investigated LLM-driven test improvements at Meta, focusing on refining correctness but without addressing key aspects such as maintainability, readability, or test smells. These studies demonstrate progress in test improvement but remain limited to correctness-focused refinements.
In contrast, our work provides a multi-dimensional assessment that goes beyond correctness refinements. We systematically evaluate test smell detection, an overlooked issue in prior work, revealing structural weaknesses in LLM-generated tests. Additionally, we incorporate execution-based correctness analysis, ensuring that refined tests not only compile but also behave correctly at runtime. Moreover, we introduce readability and maintainability metrics, assessing whether refined tests are understandable and practically usable in real-world software projects.
Our findings indicate that while iterative refinement can enhance LLM-generated tests, it does not fully address fundamental readability and execution challenges. This underscores the need for hybrid approaches, combining Search-Based Software Testing (SBST) techniques with LLM-based test generation to achieve both correctness and maintainability improvements.
6.6 Key Takeaways and Research Gap
Prior works on LLM-based test generation primarily focus on correctness and coverage, leaving key aspects unexplored. Our study fills this gap by providing a comprehensive evaluation framework that goes beyond static correctness checks. We incorporate execution-based correctness assessment, analyzing runtime behaviors through pass/fail/timeout classification and failure patterns. Additionally, we assess readability and maintainability using deep learning-based readability scoring, aligning with human evaluation.
Unlike previous studies that explore only one or two prompting techniques, we systematically compare five strategies—ZSL, FSL, CoT, ToT, and GToT. We also integrate mutation testing, a missing component in prior research, to evaluate fault detection capabilities. Beyond correctness and structural coverage (line, instruction, method), we analyze test smells, code extraction reliability (MSR & CSR), compilation errors, and runtime failures, uncovering key challenges in dependency resolution and API usage.
A key differentiator is our diverse dataset selection: SF110 (academic), Defects4J (real-world bugs), and CMD (industrial codebase). This ensures our findings generalize across different software environments. By bridging correctness, execution reliability, maintainability, dataset diversity, and fault detection, our study sets a new benchmark for evaluating LLM-generated test suites in
7 Discussion of Findings and Implications
Our study reveals that reasoning-based prompting (CoT, GToT) improves test correctness and maintainability, yet execution reliability and fault detection remain significant challenges. Many tests compile but fail to verify program behavior, while test smells affect maintainability. Hybrid approaches integrating SBST and human oversight are essential for robustness. Though prompt engineering improves reliability, hallucination-induced errors persist, indicating the need for better contextual awareness in handling dependencies and encapsulation. Execution-based validation is crucial, as static correctness alone is insufficient for real-world use. These findings emphasize the need for adaptive test generation pipelines incorporating automated validation, dependency resolution, and structured refinement.
7.1 Implications for Researchers
Our findings emphasize reasoning-based prompting’s significant role in improving LLM-generated test suites. Unlike prior studies (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024), which didn’t extensively analyze prompt engineering, we systematically evaluate ZSL, FSL, CoT, ToT, and GToT. Results show structured reasoning significantly enhances test adherence, compilability, and reliability, with GToT consistently outperforming other strategies.
Our multi-level coverage analysis reveals LLM-generated tests achieve moderate coverage but remain inferior to EvoSuite. This demonstrates the need for hybrid approaches integrating SBST techniques with LLMs. Our findings also challenge the assumption that larger models produce better tests, as GPT-3.5-Turbo occasionally outperforms GPT-4 on CMD, highlighting the importance of investigating how different architectures perform across diverse scenarios.
A critical takeaway is the oracle problem in test failure interpretation. While prior work focused on defect detection through failing tests, our analysis shows many ”Passed” tests are ineffective for fault detection, often targeting interfaces, abstract classes, or trivial assertions. This underscores the necessity of developing better validation techniques for meaningful defect identification.
Our readability and maintainability analysis reveals that while LLMs generate syntactically correct and readable tests, readability alone doesn’t ensure maintainability. Test smells—particularly Magic Number Tests and Assertion Roulette—persist across all models and prompting strategies, increasing debugging complexity. This highlights the need for prompting refinements and post-processing techniques to reduce test smells and improve long-term reliability.
7.2 Implications for Developers
Our study highlights LLM-generated test suites’ practical value when readability and maintainability are prioritized over coverage. Unlike EvoSuite, which maximizes coverage at the expense of readability, LLM-generated tests offer better structure, clearer naming, and improved comprehensibility—especially when using reasoning-based prompting like GToT and CoT. This makes these tests more accessible for long-term maintenance.
However, developers must remain aware of reliability concerns. While LLMs generate tests rapidly, compilability doesn’t guarantee execution success. Our analysis reveals many LLM-generated tests fail or time out, reinforcing the need for manual validation before pipeline integration. Although these tests achieve moderate coverage, mutation testing suggests many don’t meaningfully contribute to fault detection, often targeting trivial code portions. Developers should supplement LLM-based generation with search-based techniques to enhance fault detection.
Persistent test smells—redundant assertions, flaky tests, improper assertions—emphasize human oversight’s importance. LLMs provide a strong starting point but should be treated as assistive tools rather than autonomous solutions. Effective prompt refinement, automated filtering, and post-processing validation maximize LLM-generated tests’ utility, ensuring meaningful contributions to software quality and long-term maintainability.
8 Conclusion
This study evaluates LLMs for automated unit test generation, examining structured prompting strategies across test quality, execution reliability, maintainability, and fault detection. We analyze four LLMs and five prompting techniques to assess how prompt engineering influences test suite quality. Our findings show reasoning-based prompting, particularly GToT, improves test structure, readability, and compilability. However, execution reliability remains challenging—many tests compile but fail to verify program behavior. Mutation testing reveals tests often target trivial paths, limiting fault detection effectiveness. Hallucination-driven errors persist despite structured reasoning. While reasoning-based prompting enhances maintainability, test smells like Magic Number Tests and Assertion Roulette remain common. Coverage analysis demonstrates LLMs achieve moderate coverage but underperform EvoSuite. These findings highlight the need for hybrid approaches integrating LLM generation with search-based testing and human oversight. Our work extends beyond static correctness by incorporating execution-based classification, hallucination analysis, and mutation testing across diverse datasets. While LLMs improve test readability and structure, they still suffer from execution failures, hallucinations, and weak fault detection. Future research should explore hybrid methodologies combining LLM generation with search-based techniques, automated validation, and hallucination mitigation to enhance software quality assurance.
Declarations
Funding
This research was funded in whole, or in part, by the Luxembourg National Research Fund (FNR), grant reference AFR PhD bilateral, project reference 17185670. This work was also supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 949014). For the purpose of open access, and in fulfilment of the obligations arising from the grant agreement, the author has applied a Creative Commons Attribution 4.0 International (CC BY 4.0) license to any Author Accepted Manuscript version arising from this submission.
Conflict of Interest
The authors declare that they have no conflict of interest.
Ethical Approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Informed Consent
Not applicable.
Author Contributions
All authors contributed equally to this work. They jointly conceived the study, conducted the experiments, analyzed the results, and wrote the manuscript. All authors reviewed and approved the final version of the manuscript.
Data Availability Statement
The datasets and code used in the present study are available in our repository: https://anonymous.4open.science/r/LLM4TS-0F76/
Clinical Trial Number
Not applicable.